### v1

# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

Abstract
In June 2022, the Gaia mission released a catalog of astrometric orbital solutions for 168,065 binary systems, by far the largest such catalog to date. Unlike previous binary stars catalogs, which were heterogeneous collections of orbits from different surveys and instruments, these orbits were derived using Gaia data alone. Despite this homogeneity, the selection function is difficult to characterize because of choices made in the construction of the catalog. Understanding the catalog’s selection function is required to model and interpret its contents. We use a combination of analytic and empirical prescriptions to construct a function that computes the probability that a binary with a given set of properties would have been published in the Gaia Data Release 3 astrometric orbit catalog. We also construct a binary population synthesis model based on Moe & Di Stefano (2017) to validate our characterization of the selection function, finding good agreement with the actual Gaia NSS catalog, with the exception of the orbital eccentricity distribution. The NSS catalog suggests high-eccentricity orbits are relatively uncommon at intermediate periods $100\lesssim P_{orb}\lesssim 1000$ days. As an example application of the selection function, we estimate the Gaia DR3 detection probabilities of the star + BH binaries Gaia BH1, BH2, and BH3. We also estimate the population of Sun-like star + BH binaries in the Galaxy to be $\sim 5000$ for $400100<P_{orb}<$ 400 day, $\lesssim 2,000$ for $1000400<P_{orb}<$ 1000 day, and $\lesssim 20,000$ for $20001000<P_{orb}<$ 2000 days.
facilities: Gaiasoftware: isochrones (Morton, 2015), galaxia (Sharma et al., 2011), astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2019), healpy (Zonca et al., 2019), mwdust (Bovy et al., 2019), COSMIC (Breivik et al., 2020)
1 Introduction
Gaia is an astrometric space mission surveying over 2 billion stars in the Milky Way (Gaia Collaboration et al., 2016a). Its main mission operated from July 2014 through December 2019, and its extended mission is planned to operate through January 2025. The data products and catalogs generated from the photometry, astrometry, and spectroscopy from Gaia are shared publicly in data releases that occur every several years.
In previous Gaia data releases, all catalog parameters were produced using single-star models (Gaia Collaboration et al., 2016b, 2018, 2021a). For the first time in Data Release 3 (DR3), a catalog of non-single stars (NSS) was produced (Gaia Collaboration et al., 2023a, b). A wide variety of science involving binary and multiple stars has already been done with the NSS catalog since it was released in June 2022, including the first mass measurement of a star orbiting a detached black hole (El-Badry et al., 2023a; Chakrabarti et al., 2023); for a summary of other discoveries, see El-Badry (2024).
Performing population studies of the stellar and compact remnant binaries found in Gaia DR3 requires understanding the selection effects of the NSS catalog. The Gaia collaboration cautions that this is non-trivial to derive from their processing pipeline (Halbwachs et al., 2023). Furthermore, all Gaia data releases so far, including DR3, only provide model parameters, and not the individual photometric, spectroscopic, or astrometric measurements; without the starting input, it is not possible to exactly reproduce the processing pipeline or perform injection and recovery tests.
In this paper, we perform an empirical forward model of the Gaia NSS astrometric orbits catalog. We focus on the solutions that have an astrometric orbit (as compared to those that only have a spectroscopic orbit) as our primary goal is to understand the selection function of stellar-mass black hole + star binaries, which have thus far only been found in the astrometric orbits catalog. Our forward modeling approach uses analytic and simple prescriptions to approximate the steps in the construction of the catalog. This is a complementary approach to that of El-Badry et al. (2024), who forward model individual observations at the expense of computational cost; more discussion on the tradeoffs between the two approaches is presented in §6.5.
The rest of this paper is organized as follows. In §2 we review the processing pipeline of the Gaia NSS astrometric orbit catalog and identify the most important steps to model. In §3 we describe the binary population synthesis model. In §4 we forward model the NSS astrometric processing cascade, and the results are presented in §5. Applications, shortcomings, possible extensions of the forward model, and comparison to other work are discussed in §6. We provide a final summary and conclusions in §7.
2 Gaia astrometric non-single star processing and catalog
The astrometric NSS processing and resultant catalog were designed with the intention of prioritizing accurate binary parameters over completeness. This resulted in a series of many filtering steps to remove potentially spurious solutions, at the expense of many accurate solutions. Our goal is to understand the selection effects of this filtering process. The astrometric NSS processing described in Halbwachs et al. (2023) is summarized below, and the portions most relevant to the forward modeling are highlighted; see also Figure 1 of Halbwachs et al. (2023) for a summary.
2.1 Astrometric orbit solutions
The Gaia DR3 NSS catalog is divided into four tables depending on the type of NSS model used; each table is further subdivided by the particular sub-model used (see §2.1. “Table contents” of Gaia Collaboration et al. (2023b)). This paper is only concerned with modeling the subset of solutions that have astrometric orbits, i.e. Orbital and AstroSpectroSB1 solutions from the the nss_two_body_orbit table. We collectively refer to these as the NSS astrometric orbit solutions.
2.2 Selection of stars to be processed
An initial filtering was performed on the entire Gaia DR3 gaia_source  catalog to define the input catalog, a subset of sources that were good NSS candidates. Starting from $1.5\times 10^{9}$ sources with single-star astrometric solutions, several filtering steps were implemented to produce the input catalog of $4.1\times 10^{6}$ stars. Only stars in this input catalog were eligible to be fit with an astrometric NSS model. As described in Halbwachs et al. (2023) §2.1, the filters are as follows:
[…]
  • $\texttt{ruwe}>1.4$ (square root of the reduced $\chi^{2}$, with an empirical renormalization based on magnitude and color of source),
[…]
  • $C^{*}<1.645\sigma_{C^{*}}$ (color excess (ratio between the total flux in BP and RP, and the G-band flux (BP + RP)/G) after correcting for passbands, possibly from a contaminating neighbor).
We only model the ruwe  filter. The ipd_frac_multi_peak, ipd_gof_harmonic_amplitude, and $C^{*}$ filters are intended to remove marginally resolved wide binaries. Such binaries tend to have orbital periods $P_{orb}\gtrsim$ decades, which are much too long to receive an astrometric orbital solution; they have $\texttt{ruwe}>1.4$ because their marginally resolved nature leads to spurious epoch astrometry. The filters on apparent magnitude and visibility periods are accounted for in the $\texttt{ruwe}>1.4$ filter. As ruwe  is a measurement of the significance of deviations from a single-star model, sources that are faint or have too few visibility periods will have astrometric uncertainties too large to have a significant deviation.
2.3 Model cascade
Subsequently, the $4.1\times 10^{6}$ stars from the input catalog were fit with a variety of astrometric NSS models. The models were tried in the following order:
  • variable acceleration (Acceleration9 solution, nss_acceleration_astro table; 9 parameters),
  • constant acceleration (Acceleration7 solution, nss_acceleration_astro table; 7 parameters),
  • orbital (Orbital solution, nss_two_body_orbit table; 12 parameters11110 parameters were used for a circular or psuedo-circular orbit (see §5.2 “Pseudo-circular and circular orbital solutions” of Halbwachs et al. (2023)). 15 parameters were used if the orbital solution was combined with a spectroscopic solution (AstroSpectroSB1 solution). The astrometry-only and spectroscopic-only inferred parameters were only combined if they matched (see §5.3. “Selection of the orbital solutions” of Halbwachs et al. (2023)).).
If an acceptable solution was found, subsequent models were not tried. Note that the acceleration models precede the orbital model. Thus, if an acceptable acceleration model was found, an orbital model was not tried, even if an orbital solution would have led to a better fit. We discuss the implications of this choice in §4.3.
2.4 Orbital solution model acceptance
For sources where an orbital model fit was attempted, there were several conditions determining whether the solution was accepted and included in the final orbital solution catalog. A stringent filter was applied on the significance of the parallax solution ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$) where $\varpi$ is the parallax, $\sigma_{\varpi}$ is the parallax error, and $P_{orb}$ is the orbital period. This filter was designed to remove false positives with spurious astrometric orbits and orbital periods related to the Gaia scanning law (Holl et al., 2023), but came at the expense of many valid solutions.
There were also several other less stringent filters. A filter was applied on the significance of the astrometric orbit $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ where $a_{0}$ and $\sigma_{a_{0}}$ are the photocenter orbit semimajor axis and its uncertainty. Another filter was applied on the error of the orbital eccentricity $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$. Finally, a filter was also applied on the goodness of fit of the binary solution $F_{2}=\sqrt{9\nu/2}[(\chi^{2}/\nu)^{1/3}+(2/9\nu)-1]<25$, where $\chi^{2}$ is the sum of the squared normalized residuals and $\nu$ is the number of degrees of freedom.
We only model the filters on $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$. The filter on $\sigma_{e}$ only removes a small percentage of solutions. As described in §5.3 “Selection of the orbital solutions” of Halbwachs et al. (2023), approximately 180,000 orbital solutions remained after applying the filters $F_{2}<25$, $a_{0}/\sigma_{a_{0}}>5$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$; subsequently applying the filters on $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$ and $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{day}}$ only removed an additional 7% of solutions. Thus, the contribution of the $\sigma_{e}$ filter cannot be more than 10% of the final NSS astrometric orbit catalog [^2] Using the modeling of El-Badry et al. (2024), we find that the $\sigma_{e}$ filter only removes 2% of sources, which is 1/3 of the 7% removed by the $a_{0}/\sigma_{a_{0}}$ filter.; this is negligible given the other uncertainties in the forward model. Visual inspection of Figure 3c of Halbwachs et al. (2023) also suggests the $\sigma_{e}$ filter has minimal impact on the final constructed catalog.
We also do not model the filtering on $F_{2}$ because we assume that good orbital solutions will have good fits. $F_{2}$ is expected to follow a normal distribution with mean 0 and standard deviation 1; even if the uncertainties are poorly characterized, $F_{2}<25$ is still a 25-$\sigma$ threshold. It may also allow false positives to pass through, but is unlikely to remove real solutions that have passed the other filtering steps (with the exception of binaries in hierarchical triple systems).
2.5 Final NSS astrometric orbit catalog
The black outline in Figure 7 shows the distribution of orbital period, parallax, eccentricity, photocenter orbit semimajor axis, renormalized unit weight error, apparent magnitude, and parallax over parallax error for the final set of 168,065 NSS astrometric orbit solutions. Some of the correlations are clearly due to the filtering procedure (e.g. orbital period vs. parallax over parallax error.)
3 Binary population synthesis
We forward model a binary population and apply the filters described in §2 to attempt to reproduce the Gaia NSS astrometric orbit catalog. Note that only binary stars are modeled; single stars are not included, nor are higher-order multiple systems.
3.1 Galactic model (positions, kinematics, ages, metallicities, extinction)
[…]
We only simulate stars within 2 kpc of the Sun, as beyond that Gaia’s sensitivity to astrometric binaries in DR3 significantly drops. Note that within 2 kpc there are no bulge stars. To validate this choice, we simulate all the stars in the Galaxy to determine how many solutions are potentially lost, using the forward model described later in this paper. Only $\approx 1\%$ of solutions are beyond 2 kpc (and all solutions are within 3.5 kpc), so limiting the simulated sample to sources within 2 kpc is justified.
We compare the star counts predicted by galaxia to those in the Gaia Catalogue of Nearby Stars (GCNS; Gaia Collaboration et al., 2021b), which is $>~{}92\%$ complete down to $0.08M_{\odot}$ within 100 pc of the Sun. We find that galaxia over-predicts the number of stars within 100 pc by a factor of about 1.5 (Figure 1). galaxia predicts 0.51 million stars within 100 pc, and 1.1 billion single stars within 2 kpc. In comparison, the GCNS contains 0.36 million sources after completeness correction within 100 pc. Assuming this trend holds out to 2 kpc and that one GCNS source is equivalent to one galaxia star, we rescale the galaxia star counts by 0.65 to match the observed source counts in the GCNS. After rescaling the galaxia star counts to match the GCNS, there are 720 million galaxia stars within 2 kpc.
Extinction was calculated using the 3-D dust map mwdust (Bovy et al., 2019), which combines the dust maps of Marshall et al. (2006); Drimmel et al. (2003); Green et al. (2019). mwdust does not have a Gaia G-band filter, so we approximate it using the Gunn r-band filter (this is equivalent to using $A_{G}=2.3E(B-V))$.
Figure 1: Number of sources as a function of distance from the Sun, for the Gaia Catalogue of Nearby Stars (GCNS; black line) vs. the galaxia simulation (blue dash). The GCNS is $>92\%$ complete for stars down to $0.08M_{\odot}$ within 100 pc of the Sun; to obtain the completeness corrected number of sources, we divide the number of GCNS sources in each histogram bin by 0.92. In order to have agreement with the completeness corrected source counts in GCNS, the number of galaxia sources must be reduced by a factor of 0.65 (blue line).
3.2 Stellar population model (binary properties and evolution, IMF, synthetic photometry)
To forward model the NSS catalog, we cannot directly use the stellar population generated by galaxia since it consists of only single stars. Hence, we combine galaxia positions, metallicities, and ages with a binary star population based on Moe & Di Stefano (2017), which is a joint probability distribution in primary mass $M_{1}$, mass ratio $q=M_{2}/M_{1}$, orbital period, and eccentricity. Moe & Di Stefano (2017) infer the distribution bover $0.8<M_{1}<40M_{\odot}$, $10.1<q<$ 1, $80.15<\log P_{orb}/\textrm{day}<$ 8, and $10<e<$ 1. For reference, about 60% of Sun-like stars are single, while this decreases to nearly 0% for OB-type stars.
We use the Compact Object Synthesis and Monte Carlo Investigation Code (COSMIC; Breivik et al., 2020) to sample this joint distribution. We generate binary systems with primary masses spanning $0.08<M_{1}<20M_{\odot}$. The primary masses in the distribution of Moe & Di Stefano (2017) only go down to $0.8M_{\odot}$, so for the low mass primaries $0.08<M_{1}<0.8M_{\odot}$, it is assumed the joint probability distribution of orbital periods and eccentricities are the same as those with primary mass $M_{1}=0.8M_{\odot}$, and the binary fraction is extrapolated such that it is 0 at $M_{1}=0.08M_{\odot}$, and the mass ratio distribution is truncated such that the minimum $q$ is $0.08M_{\odot}/M_{1}$.
We assume that the mass distribution of primaries and singles are drawn from a Kroupa initial mass function (Kroupa, 2001). This results in a multiplicity frequency [^4] Following the definition of Reipurth & Zinnecker (1993), the number of binary systems divided by the number of binary systems and single stars. of 0.21 and a companion probability [^5] Again following Reipurth & Zinnecker (1993), the number of stars in binary systems divided by the total number of stars. of 0.35. This multiplicity frequency might seem significantly lower than binary frequency of $30-40\%$ typically quoted for solar-type stars (e.g., Raghavan et al., 2010; Moe & Di Stefano, 2017). This is because the binary statistics in Moe & Di Stefano (2017) only extend down to $0.8M_{\odot}$; to extrapolate the multiplicity statistics down to $0.08M_{\odot}$, COSMIC assumes the binary fraction decreases linearly from $\approx 40$% at $0.8M_{\odot}$ down to 0 at $0.08M_{\odot}$. This choice of implementation is somewhat arbitrary (given that M-dwarfs appear to have a binary frequency $\sim 20\%$; e.g., Laithwaite & Warren 2020), but ultimately these sources are too faint to be included in the astrometric orbits catalog, so their assumed binary fraction does not matter.
Assuming that the majority of binary systems are unresolved, one galaxia star corresponds to one COSMIC stellar system. This corresponds to 150 million binary systems within 2 kpc at the present day. In order to properly generate the birth population, the stellar progenitors of the compact objects must be accounted for [^6] Note that compact objects are not included in galaxia.. We use the MIST stellar evolution models (Dotter, 2016; Choi et al., 2016) to evolve a population of stars to the same age distribution of the thin disk, thick disk, stellar halo, and bulge. We then determine how many of the primary stars have or will imminently become compact objects (defined as carbon burning for massive stars, and the end of the post-AGB phase for low/intermediate mass stars), and apply a correction factor to the number of stars that need to be generated. This correction is 7.5% for the 5–7 Gyr old population in the thin disk, 11.5% for the 11 Gyr old population in the thick disk, and 12.5% for the 13 Gyr old population in the halo. The final result is a birth population of 170 million binary systems within 2 kpc.
For binaries where the primary has reached carbon burning, we assume it has evolved into a neutron star or black hole, and the binary is disrupted by the natal kick of the neutron star or black hole. This means there are no binary neutron stars nor black holes in the population synthesis. This may seem strange given that one of the goals of this work is to determine the number of black holes in wide binaries in Gaia. However, the goal of the population synthesis is not to determine how many black hole + star binaries are produced; rather, the population synthesis is used to evaluate whether or not our framework for forward modeling the NSS astrometric catalog can reproduce the actual catalog. Detailed population synthesis studies (e.g., Yamaguchi et al., 2018; Wiktorowicz et al., 2020; Chawla et al., 2022) estimate there to be $10^{5}-10^{6}$ black hole + star binaries in the entire Galaxy; within 2 kpc (which contains about 1% of Galactic stars), there will be $10^{3}-10^{4}$ black hole + star binaries. After applying detectability criteria, the previously mentioned studies estimate $10^{1}-10^{3}$ black hole + star binaries detectable by Gaia; these are optimistic estimates as these are end-of-mission (i.e., DR4). Thus, black hole + star binaries are a negligible part of the forward model, considering there are $O(10^{5})$ astrometric solutions total.
For binaries where both the primary and secondary have reached the end of the post-AGB phase, we assume those are two white dwarfs that will not produce a Gaia-detectable binary as both components are too faint. For binaries where only the primary has reached the post-AGB phase, we assume these are white dwarf + star binaries. Following the prescription of El-Badry et al. (2024); Yamaguchi et al. (2024), we assume that 10% of these white dwarf + star binaries with initial separations between $2-6$ AU form wide post-common envelope binaries with separations that are half their initial separation. The white dwarfs are assumed to be dark and their masses are determined from their zero-age main sequence mass using the initial-final mass relation of Kalirai et al. (2008). The remaining white dwarf + star binaries are assumed to either have orbital periods too short ($\lesssim 10$ days) or too long ($\gtrsim 10^{4}$ days) that they could not be detected astrometrically by Gaia in DR3. We also remove binaries where either star is currently or has previously overflowed its Roche lobe, under the assumption that the periods will be too short to detect astrometrically in Gaia DR3 due to orbital shrinking.
Because of the uncertainties in the effects of interactions in binary stars and the precise properties of natal kicks in neutron stars and black holes, we do not evolve the binary systems using COSMIC; we only use COSMIC to generate the population of binary stars in their birth configuration. The majority of binaries in the NSS astrometric catalog are main sequence + main sequence binaries that have never interacted, so treatment of evolutionary effects is less important than the modeling of the Gaia astrometric processing cascade.
The MIST models are also used to generate synthetic Gaia G-band photometry for the stars as a function of mass, age, and metallicity. Again, no binary evolution or interaction is taken into account; each star is individually evolved using a single-star isochrone.
With the population simulated, we can calculate any other properties of the binaries needed. Most important for this forward model is the angular semimajor axis of the orbit photocenter
$a_{0}=a\varpi\delta_{ql},$ (1)
where $a$ is the semimajor axis of the binary (in physical units), $\varpi$ is the parallax, and $\delta_{ql}$ is the photocenter scale. The photocenter scale, the factor by which the photocenter orbit semimajor axis is smaller than the semimajor axis, is defined as
$\delta_{ql}\equiv\Bigg{|}\frac{q_{l}}{q_{l}+1}-\frac{l}{1+l}\Bigg{|}$ (2)
where $l$ is the luminosity ratio $l=L_{2}/L_{1}$ of the binary system and $q_{l}$ is the mass ratio $q_{l}=M_{2}/M_{1}$ (van de Kamp, 1975; Belokurov et al., 2020). By definition here, the primary (component 1) is brighter, but not necessarily more massive. The mass ratio $q$ is defined to be $0\leq q\leq 1$. Thus, $q_{l}=q$ if component 1 is both brighter and more massive, otherwise $q_{l}=1/q$. Note that the photocenter orbit semimajor axis is maximized for dark companions (i.e., $l=0$) and large mass ratios (i.e., $q_{l}\gg 1$), and is zero for equal-mass companions of the same luminosity ($l=1$, $q_{l}=1$). Finally, we note that all these luminosity-dependent quantities are specifically in the Gaia G-band.
4 Empirical modeling of Gaia NSS model cascade
With a binary population model in hand, the next step is to reproduce the modeling cascade described in §2. In this section, we describe how we estimate and implement the filters $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$. In addition, we also need a way to account for the fact that 1) Gaia DR3 is based on only 34 months of data, 2) orbital solutions are “lost” higher up in the model cascade when they are fit with an acceleration solution, and 3) orbital periods of 1 year are very difficult to distinguish from parallax. We describe the empirical filters we apply to account for these effects.
4.1 Empirical ruwe  model
The renormalized unit weight error ruwe  is used to quantify the goodness-of-fit of a single star model. Specifically, it is the square root of the reduced $\chi^{2}_{r}$ (the “uwe” in ruwe) with a correction applied to mitigate systematics in the $\chi^{2}$ distribution due to color and magnitude (the “r” in ruwe; Lindegren 2018).
4.1.1 Analytic ruwe  prediction
Belokurov et al. (2020) demonstrated a close relationship between ruwe  and the presence of non-single stars using Gaia DR2 data (which did not include any type of binary or NSS models). If $\delta\theta$ is the astrometric excess noise due to fitting an unresolved binary with a single star model, then
$\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1},$ (3)
where $\sigma_{AL}(G)$ is the along-scan centroiding error, which is a function of the source magnitude $G$.
Here, as well as in the remainder of this paper, we adopt the per-CCD centroiding error from Figure 3 of Holl et al. (2023). As described in Lindegren (2018), for $12\lesssim G\lesssim 17$ the formal centroiding precision is limited by photon noise (leading to Poisson errors), for $G\gtrsim 17$ background noise dominates and the error rises more quickly than Poisson, and for $G\lesssim 12$ the CCD gates limit the number of photons and the centroiding precision is independent of magnitude. The $\sigma_{AL}(G)$ we adopt from Holl et al. (2023) is empirically determined from Gaia DR3 data, and is significantly larger ($>0.3$ dex) than the formal error for bright ($G\gtrsim 12)$ sources.
[…]
Penoyre et al. (2020) derived an analytic relationship (their Equation 17) between the astrometric excess noise $\delta\theta$, the source magnitude $G$, photocenter orbit semimajor axis $a_{0}$, inclination $i$, and eccentricity $e$
$\delta\theta=a_{0}\sqrt{1-\frac{\sin^{2}i}{2}-\frac{3+\sin^{2}i(\cos^{2}\phi_{\nu}-2)}{4}e^{2}},$ (4)
where $\phi_{\nu}$ is the azimuthal viewing angle of the binary. Note that the derivation of Equation 4 assumes that there are many observations sampling the full orbital period of the binary during the survey, such that the observations can be time-averaged over the orbit.
For an isotropic distribution of $\phi_{\nu}$, the average value of $\cos^{2}\phi_{\nu}$ is 0.5. Fixing $\cos^{2}\phi_{\nu}$ to 0.5, the astrometric excess noise can be re-written
$\delta\theta=a_{0}\sqrt{\frac{1}{2}\left(1+\cos^{2}i\right)\left(1-\frac{3e^{2}}{4}\right)}.$ (5)
Combining Equations 3 and 5 results in a formula for ruwe  as a function of source magnitude, photocenter orbit semimajor axis, eccentricity, and inclination:
$\texttt{ruwe}=\sqrt{1+\frac{1}{2}\left(\frac{a_{0}}{\sigma_{AL}(G)}\right)^{2}\left(1+\cos^{2}i\right)\left(1-\frac{3}{4}e^{2}\right)}.$ (6)
Inspection of Equation 6 shows that face-on orbits ($|\cos i|\approx 1$) have larger ruwe  than edge-on orbits ($|\cos i|\approx 0$), and circular orbits ($e\approx 0$) have larger ruwe  than highly eccentric orbits ($e\approx 1$). If the orbit is perfectly face on and circular ($i=0$, $e=0$), this suggests that the astrometric excess noise is equivalent to the semimajor axis of the photocenter orbit (note that $\delta\theta\leq a_{0}$). Thus, on average, filtering on a minimum value of ruwe  favors circular and face-on orbits. However, ruwe  depends more strongly on inclination than eccentricity: the term containing inclination $1+\cos^{2}i$ spans the interval [1, 2], but the term containing eccentricity $41-3e^{2}/$ 4 only spans the interval [0.25, 1]. We also visually validate the scaling between the astrometric excess noise and photocenter orbit semimajor axis, inclination, and eccentricity in Figure 2.
Figure 2: Using the NSS astrometric orbits catalog, we validate the scaling relations between the approximation of the astrometric excess noise $\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1}$ and the semimajor axis of the photocenter orbit $a_{0}$, eccentricity $e$, and orbital inclination $i$. The left column shows the approximation of $\delta\theta$ vs. $a_{0}$, with the color of the point denoting $|\cos i|$ (top panel) or eccentricity of the orbit (bottom panel). There is a linear trend between $a_{0}$ and $\delta\theta$ where the slope is less than unity, as expected. In addition, there is a clear gradient in the colors of the points: averaging over the orbit, face-on and circular orbits $(|\cos i|\approx 1,e\approx 0)$ have larger astrometric excess noise at fixed $a_{0}$ than edge-on and highly eccentric orbits $(|\cos i|\approx 0,e\approx 1)$. The right column is the same as the left column, but includes the dependence of $e$ and $i$ on $\delta\theta$, i.e. shows the approximation of $\delta\theta$ vs. $a_{0}\sqrt{(1+\cos^{2}i)(1-3e^{2}/4)}$. The trends in the colors of the points mostly disappear (the $|\cos i|$ and $e$ dependence), validating the scaling relation. There are still some minor trends visible in $|\cos i|$, but Gaia Collaboration et al. (2023b) showed that the sensitivity of circular orbits deviates slightly from the the expected $1+\cos^{2}i$, so this is not surprising. Note that there are very few red points in the bottom panels because there are very few high-eccentricity orbits.
4.1.2 Empirical correction to analytic ruwe  prediction
The predicted astrometric excess noise in Equation 5 and ruwe  in Equation 6 will generally be overestimates of the astrometric excess noise and ruwe  reported by Gaia. The derivation of Equation 5 assumes that the binary signal has not been erroneously absorbed into other astrometric terms (i.e., position, parallax, proper motion, or satellite pointing noise). Penoyre et al. (2020) predicted that this erroneous absorption would happen frequently, often lowering the reported astrometric excess noise and ruwe. With the NSS catalog, we confirm that the predicted astrometric excess noise $\delta\theta$ tends to be smaller than the actual $\delta\theta$ in the catalog by $\approx 60-70\%$ (Figure 3).
We calculate an empirical correction to rescale the analytic ruwe  prediction to match the actual ruwe  using the NSS catalog. The number of visibility periods used in the solution ($n_{vis}$) affects the size of the erroneous absorption. For smaller $n_{vis}$, the average size of the erroneous absorption is larger, and there is more scatter about this average; for larger $n_{vis}$, the average size of the erroneous absorption is smaller, and there is less scatter about this average (Figure 3).
To calculate the correction factor, we split the NSS astrometric orbit solutions into three bins based on the number of visibility periods used: $n_{vis}<23$, $2523\leq n_{vis}<$ 25, and $n_{vis}\geq 25$. These bins were chosen since $n_{vis}$ is a bimodal distribution for the NSS astrometric orbits catalog, with peaks at $n_{vis}=18$ and 27. In each of these bins, a Gaussian is fit to the distribution of the ratio of the actual $\delta\theta$ to the analytically predicted $\delta\theta$ from Equation 5. The mean $\mu$ and standard deviation $\sigma$ for each of the bins are $\mu=0.574$, $\sigma=0.133$ for $n_{vis}<23$; $\mu=0.606$, $\sigma=0.116$ for $2523\leq n_{vis}<$ 25, and $\mu=0.643$, $\sigma=0.099$ for $n_{vis}\geq 25$. To apply this correction, a sample is drawn from the appropriate Gaussian and used to rescale the analytic estimate $\delta\theta$, and in turn ruwe, since $\texttt{ruwe}=\sqrt{1+(\delta\theta/\sigma_{AL})^{2}}$ (Equations 4 and 5). We note that we chose not to directly calculate the correction using the ruwe  reported in the NSS astrometric catalog, because the hard limit $\texttt{ruwe}>1.4$ makes the distribution non-Gaussian and would be more difficult to model than using $\delta\theta$.
Figure 3: Comparison of the astrometric excess noise $\delta\theta$ for sources in the NSS astrometric orbit catalog, vs. the analytic predicted $\delta\theta$ of those sources calculated using the photocenter orbit semimajor axis, eccentricity, and inclination from the NSS catalog and Equation 5. We perform the comparison by the number of visibility periods used $n_{vis}$ in the solution. The top panels show scatter plots of the actual $\delta\theta$ vs. analytic predicted $\delta\theta$, with the 1:1 relation is denoted by the orange line. The bottom panels show the ratio of the actual to analytic predicted $\delta\theta$, with the best-fit Gaussian shown in the black line. The predicted astrometric excess noise is systematically higher than the actual value reported in the NSS catalog. The actual values reported are typically only $\approx 60-70$% of the predicted values and do not depend on the size of $\delta\theta$; the values are more severely underestimated when smaller numbers of visibility periods are used, and the scatter about the average is also larger. Some of the astrometric excess noise being erroneously absorbed into other parts of the astrometric model is likely one contributing factor to the underestimation of $\delta\theta$.
4.2 Empirical parallax and photocenter orbit semimajor axis error
We construct an empirical error model for the parallax and photocenter orbit semimajor axis using the NSS catalog.
The parallax error $\sigma_{\varpi}$ and photocenter orbit semimajor axis $\sigma_{a_{0}}$ are expected to be related to the centroiding error in the along-scan direction, $\sigma_{AL}$, and the number of visibility periods used, $N_{vis}$, according to
$\sigma_{\varpi},\sigma_{a_{0}}\propto\frac{\sigma_{AL}(G)}{\sqrt{N_{vis}}}.$ (7)
Using the number of visibility periods provides a more reliable estimate for position and proper motion uncertainties than the total number of good along-scan observations (Lindegren et al., 2018).
To obtain an estimate for $N_{vis}$ in the population synthesis, we use the Gaia Observation Forecast Tool (GOST) [^7] https://gaia.esac.esa.int/gost/, which predicts when any given part of the sky crosses Gaia’s focal plane. To calculate $N_{vis}$, we evaluate the number of predicted scans at 49152 evenly spaced points on the sky, discard a random 10% of the predictions to mimic losses due to dead time [^8] The GOST page suggests 20% would be more appropriate, but we find 10% gives good agreement with observations., then calculate how many of the remaining predicted observations are separated by at least 4 days.
We first attempted to find the constant of proportionality in Equation 7 relating $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ to $\sigma_{AL}(G)/\sqrt{N_{vis}}$. However, we find there is likely an additional unaccounted-for variable in the magnitude of the source (Appendix A).
Thus, we instead rely on the uncertainties from the NSS catalog to make the estimates of $\sigma_{\varpi}$ and $\sigma_{a_{0}}$. In bins of 0.2 mag and integer values of visibility periods, we calculate the median and standard deviation $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in the NSS catalog. As expected, having more visibility periods and/or having a source magnitude $8\lesssim G\lesssim 13$ (where $\sigma_{AL}$ is smallest) tends to result in smaller values of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ (Figure 4). We find the standard deviation is roughly constant when the bin is well-sampled; we thus adopt a constant standard deviation of 0.12 mas for $\log_{10}\sigma_{\varpi}$ and 0.37 mas for $\log_{10}\sigma_{a_{0}}$. We also calculate the minimum value of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in each bin.
Figure 4: Median parallax error (left) and photocenter orbit semimajor axis error (right) interpolated using the NSS catalog. The black dots show the location of the grid points in NSS; the inferred interpolated values are shown as the colored boxes. Note the color bar scale is different in the two panels.
The median, standard deviation, and minimum values are used as the mean $\mu$, standard deviation $\sigma$, and minimum $a$, respectively, in a truncated normal distribution
$f(x;\mu,\sigma,a,b)=\begin{dcases}\frac{1}{\sigma}\frac{\phi(x;\mu,\sigma)}{\Phi(a;\mu,\sigma)-\Phi(b;\mu,\sigma)},&a\leq x\leq b\\ 0,&\text{otherwise},\end{dcases}$ (8)
where $\phi$ and $\Phi$ are the distributions of a normal PDF and a normal CDF, respectively. The truncation on the low end prevents artificially low errors; we do not apply the truncation on the higher end (hence the parameter $b$ in Equation 8 is infinite; in practice, it is set as a large number like 999).
We then interpolate across this grid of values to determine the median $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ for a given magnitude and number of visibility periods used. Note that because this interpolation is performed over the NSS catalog, the interpolation is not defined for all pairs of ($N_{vis}$, $G$); the relationship would need to be extrapolated in order to estimate $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ at fainter magnitudes.
The main concern with this approach is that by using the NSS catalog itself to estimate the errors, this will propagate whatever selection effects we are trying to characterize into the prediction itself. However, since the final forward model provides reasonable agreement with observations, we conclude that this approach is acceptable, if not ideal.
4.3 Lack of sensitivity at $P\approx 1$ year and lack of solutions with $P>550$ days
Figure 5: Empirical function that removes orbital periods around 1 year and at long orbital periods $P_{orb}\gtrsim 550$ days. This empirical filter is applied after the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters. The data used to make this figure are available through “Data behind the Figure”.
The black outline in the topmost 1-D histogram in Figure 7 shows the distribution of all NSS astrometric orbital periods in Gaia. There is a clear dearth of solutions around 1 year because the astrometric signature of annual parallax is degenerate with $P_{orb}=1$ year. The 1-D histogram also shows a steep drop-off in the number of solutions at longer orbital periods. This is expected for $P_{orb}\gtrsim 1000$ days due to Gaia DR3 only containing 34 months of data. However, the number of solutions rapidly drops around $P_{orb}\approx 550$ days, which is much shorter than the Gaia DR3 temporal baseline.
Forward modeling indicates that the ruwe  and parallax over parallax error filters do not cause the peak of the period distribution to be less than 2000 days (§5.1.2). We hypothesize that the drop-off in solutions at 550 days is primarily due to the astrometric processing cascade. As described in §2.3, acceleration solutions are tried before orbital solutions in the model cascade, and if an acceptable acceleration solution is found, it is kept and an orbital solution is not searched for, even if an orbital solution would have resulted in a better fit. Thus, long-period binaries are more likely to be satisfactorily fit with an acceleration model than binaries with orbital periods short enough to be covered multiple times over DR3.
We test this hypothesis by cross-matching the nss_acceleration_astro  catalog to external binary catalogs to empirically estimate how many orbital solutions are lost to acceleration solutions in the model cascade (Appendix B). After accounting for the number of solutions at a given orbital period, we confirm that the number of acceleration solutions increases at longer periods, and the number of orbital solutions decreases at longer periods. This is also confirmed by the modeling by El-Badry et al. (2024) (their Figure 11).
Modeling the dearth of solutions at $P_{orb}\approx 1$ year and the decrease in number of long $P_{orb}$ solutions require detailed knowledge of the spacecraft scanning law and individual epoch data. In the spirit of keeping calculations fast and analytic in this work, we construct a probabilistic function to empirically account for these effects. For $P_{orb}<250$ days, all solutions are kept; we assume that the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters fully capture the astrometric cascade. [^9] This is not entirely accurate; examination of the period distribution show that solutions are clearly lost around 190 days, which is 3×3$\times$3 × Gaia’s precession period. However, the number of solutions is so small we ignore this. We then assume the remaining discrepancy between the Gaia NSS astrometric orbits catalog and the forward model period distributions are due to the scanning law, finite Gaia DR3 temporal baseline, and loss of orbital solutions to acceleration solutions. We empirically determine this function by dividing the number of NSS orbital solutions by the number of solutions in the forward model after application of the ruwe  and $\varpi/\sigma_{\varpi}$ filters. We normalize the distribution assuming no solutions are lost at $P_{orb}=250$ days.
Figure 5 shows the probabilistic function as a function of orbital period. We do not attempt to quantitatively understand the particular form of this function as this requires epoch-level modeling, which is outside the scope of this work. We also note that this empirical function is a function of $P_{orb}$ alone, which again is not completely accurate; for example, the 1 year orbit degeneracy can be more easily broken by highly eccentric orbits than circular orbits.
5 Results
5.1 Mock NSS astrometric orbits catalog
We investigate the effects of the filtering steps on the number of solutions and the distributions of orbital and binary properties. We also present the resulting parameter distribution of the forward model.
5.1.1 Number of solutions
Table 1 summarizes the effect of the filters on the number of orbital solutions. The forward model begins with 150 million binaries within 2 kpc of the Sun at the present day. Even before applying the filters discussed in the previous section, many binaries are simply not detectable because they are too faint ($G\lesssim 21$) to be detected by Gaia, or their orbital periods are very long ($P_{orb}\gtrsim$ 30 years). In order to generate a more realistic estimate on the number of potentially detectable Gaia binaries, we calculate the number of binaries with $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which represents the range of orbital periods and magnitudes spanned in the NSS astrometric orbits catalog; 13 million binaries satisfy these criteria. Ultimately, after applying all the filtering cuts, we find $3.5\times 10^{5}$ binaries will be included in the NSS catalog.
This is double the number of sources actually included in the NSS astrometric orbits catalog (168,065). There are several possible explanations for this overestimate. One possibility is that the initial binary population produced by the binary population synthesis was incorrect in some way, e.g., the absolute number of systems was too high, the orbital period distribution of the binaries were incorrect and produced too many systems with $P_{orb}\approx 100-1000$ days, or the extinction was underestimated. Another is that the filtering steps do not accurately represent the model cascade, e.g., estimating the uncertainties $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ using the NSS astrometric catalog itself biases their values low. Nonetheless, the forward model is able to reproduce the total number of sources within a factor of two, which will still provide useful constraints in population studies.
Table 1: NSS filters and resultant number of predicted binaries
Filter No. of binaries (%)
None (all binaries) $1.5~{}\times~{}10^{8}$ (100%)
$15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag $1.3~{}\times~{}10^{7}$ (8.8%)
ruwe$>1.4$ $6.3~{}\times~{}10^{6}$ (4.2%)
$\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ $1.2~{}\times~{}10^{7}$ (8.2%)
$a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{ day}})$ $8.6~{}\times~{}10^{6}$ (5.8%)
ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ $5.6~{}\times~{}10^{6}$ (3.7%)
ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$, $P_{orb}\approx$ 1 yr, $P_{orb}>550$ day $3.5~{}\times~{}10^{5}$ (0.2%)
Note. — The first line, where Filter is “None”, denotes the number of binaries within 2 kpc of the Sun. The filter on $15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag is not an actual filter criterion for the NSS astrometric orbits catalog, but is a more representative number of the possible number of orbital systems that could be detected by Gaia as compared to all binaries within 2 kpc, which has many systems with orbital periods too long or magnitudes too faint to be detected.
5.1.2 Distribution of binary and orbital parameters
Figure 6: Effect of filtering on $\texttt{ruwe}>1.4$ (ruwe), $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ ($\varpi/\sigma_{\varpi}$), $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ ($a_{0}/\sigma_{a_{0}})$, and $P_{orb}\approx 1$ yr and $P_{orb}>550$ day ($P$) on the period distribution in the forward modeling.
Figure 7: 1-D and 2-D PDFs in the period $P_{orb}$, parallax $\varpi$, eccentricity $e$, photocenter orbit semimajor axis $a_{0}$, ruwe, apparent magnitude $G$, and parallax over parallax error $\varpi/\sigma_{\varpi}$ of the forward model (red) as compared to the actual NSS astrometric orbits catalog (black). In general, the forward modeling does a good to reasonable job in matching the actual NSS distributions including the correlations, with the exception of the eccentricity distribution.
Figure 8: Effect of filtering on the distribution of orbital period $P_{orb}$, parallax $\varpi$, eccentricity $e$, ruwe, magnitude $G$, primary and secondary mass $M_{1},M_{2}$, luminosity ratio $l_{G}$, mass ratio $q$, and photocenter scale $\delta_{ql}$. The dotted line shows the birth distribution from Moe & Di Stefano (2017). The black outline shows the distribution of systems that have orbital periods between $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia. The filter on long-period orbits preferentially removes larger photocenters and large ruwe. The magnitude cut preferentially removes distant and low-mass binaries. It causes the primary mass distribution to peak roughly around $0.8M_{\odot}$ (when binned logarithmically). The gray outline shows the systems that satisfy $\texttt{ruwe}>1.4$. Finally, the red outline shows the systems that satisfy all the filtering criteria applied ($\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$, $P_{orb}\approx 1$ year, $P_{orb}>550$ day).
Figure 6 summarizes the effect of the filtering steps on the period distribution. The $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ filters all preferentially remove shorter period systems. The ruwe  filter removes more systems overall, while $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$ filters remove shorter period solutions more strongly.
Figure 7 is a corner plot comparing the orbital period, parallax, eccentricity, photocenter orbit semimajor axis, ruwe, apparent magnitude, and parallax over parallax error of the mock forward model to the actual NSS astrometric orbit catalog. The 2-D correlations are broadly preserved, and the 1-D distributions are also generally reasonable matches. The deficit of solutions around $P_{orb}\approx 1$ year and the decline in number of solutions longer than $P_{orb}>550$ days in the forward modeled period distribution were tuned to match the NSS catalog, so they are unsurprisingly a reasonable match. The parallax, magnitude, and parallax over parallax error distributions are also a very good match. The photocenter orbit semimajor axis and ruwe  distributions are also reasonable matches, although the models tend to result in distributions slightly more sharply peaked than the NSS catalog. The eccentricity distribution between the model and the true catalog are by far the most discrepant, with the Gaia NSS distribution peaking around $e\approx 0.2$, while the forward modeled distribution peaks at a much higher eccentricity of $e\approx 0.75$. We discuss possible reasons for this in §6.1. Although the parallax error and photocenter orbit semimajor axis errors were interpolated directly using the NSS catalog itself, the filters and distributions involving $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ were not tuned post-facto like the period distribution.
Figure 8 compares the 1-D distribution of binary parameters in the initial birth population with that obtained from a simple filter on observability, which we define to be $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia. It also shows the distribution of parameters after application of the $\texttt{ruwe}>1.4$ filter and all NSS filters. As discussed previously, the ruwe  filter trims short orbital period, small $a_{0}$, distant, and faint systems. Including the rest of the filters does not significantly change the 1-D distribution of the parameters as compared to the $\texttt{ruwe}>1.4$ filter, except the orbital period distribution.
5.2 Probability of finding Gaia black holes
Table 2: Gaia black hole properties and detection probabilities
Name RA (deg) Dec (deg) $P_{orb}$ (day) $e$ $\varpi$ (mas) $M_{*}(M_{\odot})$ $M_{\bullet}(M_{\odot})$ $G_{*}$ (mag) $M_{G,0*}$ (mag) $A_{G}$ (mag) p(NSS)
Gaia BH1 262.17 –0.58 186 0.45 2.09 0.93 9.62 13.77 4.6 0.76 0.33
Gaia BH2 207.57 –59.24 1277 0.52 0.859 1.07 8.94 12.28 1.8 0.15 0.04
Gaia BH3 294.83 14.93 4195 0.73 1.67 0.76 32.7 11.23 2.12 0.22 0.00
Note. — Properties of Gaia BH1, BH2, and BH3. * subscript denotes properties of the stellar companion, $\bullet$ subscript denotes properties of the black hole. All values except $M_{G,*}$ and $A_{G}$ are taken from El-Badry et al. (2023a), El-Badry et al. (2023b), and Gaia Collaboration et al. (2024) for Gaia BH1, BH2, and BH3, respectively. We note for Gaia BH2 that our extinction map suggests much less extinction ($\approx 0.4$ mag) than the map used by El-Badry et al. (2023b) and so our inferred extinction-corrected absolute magnitude of the star $M_{G,0*}$ in Gaia BH2 is fainter. Similarly for Gaia BH3 our extinction map suggests $\approx 0.3$ mag less extinction than Gaia Collaboration et al. (2024).
We apply our forward modeling framework to determine the probability of inclusion of Gaia BH1 (El-Badry et al., 2023a; Chakrabarti et al., 2023), Gaia BH2 (El-Badry et al., 2023b), and Gaia BH3 (Gaia Collaboration et al., 2024) in the Gaia DR3 NSS astrometric orbits catalog. We calculate the probability as a function of sky position, orbital period, eccentricity, parallax, and the component magnitudes and masses. We integrate over random inclinations in $|\cos i|$.
We find the probability of Gaia BH1 ((RA, Dec) = (262.17, –0.58) deg, $P_{orb}=186$ day, $e=0.45$, $\varpi=2.09$ mas, $M_{*}=0.93M_{\odot}$, $M_{\bullet}=9.62M_{\odot}$, $G_{*}=13.77$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 33%. We also vary the location of BH1 on the sky and its distance from the Sun (Figure 9) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog changes. Since Gaia BH1 was located near the ecliptic plane, it had a relatively low probability of inclusion; toward the ecliptic poles, the inclusion probability would have been closer to 70–90%. Gaia BH1 was also relatively close by, at a distance of about 500 pc. Had it been at the same (RA, Dec) but slightly farther away, at 600 pc or more, it would not have made it into the NSS catalog. Conversely, had it been slightly closer, 400 pc or less, it would have had $>80\%$ probability of making it into the catalog. This suggests that if there were Gaia BH1-like systems within a few hundred pc of the Sun, they should have been included in the Gaia DR3 NSS astrometric orbits catalog; the lack of such detections suggests that there are no or few Gaia BH1-like sources within 300–400 pc of the Sun.
We find the probability of Gaia BH2 ((RA, Dec) = (207.57, –59.24), $P_{orb}=1277$ day, $e=0.52$, $\varpi=0.86$ mas, $M_{*}=1.07M_{\odot}$, $M_{\bullet}=8.94M_{\odot}$, $G_{*}=12.28$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 4%. We also vary the location of BH2 on the sky and its distance from the Sun (Figure 9) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog varies. Since Gaia BH2 has a long orbital period, the main factor of whether it made it into the NSS catalog is whether or not it made it through the astrometric cascade or not. The inclusion probability of Gaia BH2 does not change significantly with distance until beyond $\gtrsim 2.5$ kpc, at which point the inclusion probability substantially drops. The inclusion probability also does not vary significantly across the sky. This suggests that the main factor of whether Gaia BH2 made it through the cascade or not is the long period filtering. The probability of an orbital solution being kept at $P_{orb}=1277$ days is 4% (Figure 5). This suggests that Gaia BH2 was a relatively lucky detection and that there are many more such systems that will be detectable in Gaia DR4, either in an updated NSS catalog, or in the epoch astrometry.
We find the probability of Gaia BH3 to be included in the Gaia DR3 NSS astrometric catalog to be 0%. This is because of the orbital period being $>4000$ days; none of these systems pass the long period filtering. This is not entirely correct, as $<0.5\%$ of solutions in the NSS astrometric orbits catalog have $P_{orb}>2000$ days, and $<0.05\%$ of solutions have $P_{orb}>4000$ days, so it is possible for such long solutions to be in the catalog. However, within a margin of a few percent, this is probably accurate. In addition, Gaia BH3 was not included in any of the NSS catalogs, orbital (nss_two_body_orbit), nor acceleration (nss_acceleration_astro, nss_non_linear_spectro).
Figure 9: Top panels: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of sky position (192 equally spaced points on sky); the color of the circle denotes the probability of inclusion. Note that the color bar scales in the two panels are different. We fix all parameters as listed in Table 2, and only vary the sky position. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The positions at which Gaia BH1 and BH2 were detected are marked as the red stars. For Gaia BH1, the detectability as a function of sky position clearly shows the pattern of the scanning law. For Gaia BH2, the detectability does not have a dependence on the sky position. Bottom panel: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of the system’s distance. We fix all parameters as listed in Table 2, and only vary the distance of the system. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The distances at which Gaia BH1 and Gaia BH2 were detected at are denoted by the dotted black lines.
5.3 Inferences on the population of Galactic detached black holes
Figure 10: Parallax vs. orbital period and black hole mass for a simulated population of detached Galactic BHs within 2 kpc (top panels: Sun-like stars + BHs, bottom panels: red giants + BHs). The properties of Gaia BH1 and Gaia BH2 are marked in the top and bottom panels, respectively. The detectability as a function of orbital period resembles the NSS period distribution. For Sun-like stars + BHs, most are detectable with $\varpi>1$ mas, but red giants + BHs can be detected with $\varpi>0.5$ mas and lower. Detectability does not strongly depend on the BH mass.
As an application of the forward model and selection function, we calculate the detection probabilities of systems like Gaia BH1 and BH2 in the Milky Way.
First, we generate a synthetic population of detached BH + star binaries. We assume the distribution of black hole masses follows the primary mass distribution inferred by gravitational wave measurements. We use the “Power law + peak” model (Talbot & Thrane, 2018), a normalized power-law distribution with negative spectral index (i.e., more low-mass BHs than high-mass BHs) with an additional Gaussian peak. The parameters for the model are inferred from the 2nd Gravitational Wave Transient Catalog (GWTC-2; Abbott et al., 2021), resulting in a mass distribution with a primary peak at $8M_{\odot}$, with a smaller excess at $33M_{\odot}$, notably consistent with the masses of the three Gaia BHs. The orbital period distribution is assumed to be logarithmically flat (e.g., Marín Pina et al., 2024). The eccentricity distribution is also assumed to be flat. From galaxia we sample the positions (RA, Dec, parallax) of Sun-like stars for Gaia BH1, and red giants for Gaia BH2. We use the mwdust 3-D dust map as described earlier to apply extinction.
For Gaia BH1-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+4.6$. We consider orbital periods $200010<P_{orb}<$ 2000 days, spanning roughly the range of detected NSS astrometric orbits. We will refer to these as Sun-like star + BH binaries, or Sun+BH binaries for short.
For Gaia BH2-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+1.8$. We consider orbital periods $2000300<P_{orb}<$ 2000 days. The minimum orbital period is 300 days such that the semimajor axis of the orbit is larger than the typical red giant radius. We will refer to these as red giant + BH binaries, or RG+BH binaries for short.
For the Sun+BH and RG+BH sources, we randomly pair positions and parallaxes with orbital periods, eccentricities, and black hole masses. Thus, there is no correlation between orbital period, parallax, BH mass, or position. We randomly sample 10 million of these pairs, then use the selection function to determine whether or not each object would have been included in the Gaia DR3 NSS astrometric orbits catalog.
Figure 10 shows the orbital periods, parallaxes, and black hole masses for the full population vs. detected systems. On average, 3.5% of Sun+BH binaries with orbital periods $200010<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable, and 16% of RG+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable [^10] On average, 7.0% of Sun+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable.. The period distribution unsurprisingly looks similar to the NSS astrometric orbits catalog. Sun+BH systems are generally difficult to detect beyond 1 kpc. RG+BH systems can be detected to at least 2 kpc because the red giant is much more luminous than a Sun-like star despite having the same mass. There is no strong dependence of the detectability of the system on black hole mass. The size of the photocenter orbit semimajor axis $a_{0}$ (Equation 1) increases by a factor of only 2.4 for a $5M_{\odot}$ vs. $50M_{\odot}$ black hole and $1M_{\odot}$ star, even though the black hole mass increases by a factor of 10.
We can use this model to make a simple argument on the occurrence rate of detached black holes in the Milky Way in various orbital period ranges. We first consider Sun+BH binaries with orbital periods $400100<P_{orb}<$ 400 days, $1000400<P_{orb}<$ 1000 days, and $20001000<P_{orb}<$ 2000 days, and separately RG+BH binaries with orbital periods $1000400<P_{orb}<$ 1000 days and $20001000<P_{orb}<$ 2000 days. We assume that Gaia BH1 (Sun+BH with an orbital period $400100<P_{orb}<$ 400 days) and BH2 (RG+BH with an orbital period $20001000<P_{orb}<$ 2000 days) are the only two star + BH binaries in the Gaia DR3 astrometric orbits catalog. We then use our model to calculate the number of detectable detached BHs within 1 kpc of the Sun within these period ranges. For the period ranges where no BH has been detected, we take this to be the upper limit on the number of those BH binaries. Assuming there are $\approx$100-200 billion stars in the Galaxy, 160 million, or 0.1% of those stars are within 1 kpc of the Sun; to convert our number of objects to within 1 kpc to the number of objects in the Galaxy, we multiply by 1000.
Table 3 lists the detection probabilities. We find that the Galaxy hosts $5000\pm 2000$ Sun+BH binaries with $400100<P_{orb}<$ 400 days, and $20,000\pm 4000$ RG+BH binaries with $20001000<P_{orb}<$ 2000 days. We also find that there are at most 2000 Sun+BH binaries with $1000400<P_{orb}<$ 1000 day, 20,000 Sun+BH binaries with $20001000<P_{orb}<$ 2000 day, and 20,000 RG+BH binaries with $20001000<P_{orb}<$ 2000 day. Note the limits and constraints are weaker for the long-period end, due to the sharp decrease in number of solutions beyond 1000 days.
Although the previous paragraphs treats the constraints on Sun+BHs separately from RG+BHs separately, in reality these constraints are related, as Sun-like stars will eventually evolve into red giants. Because the evolutionary timescale of the red giant phase is only a $\sim 10\%$ of a star’s main sequence lifetime, the number of RG+BHs should also be a few percent of the number of Sun+BH binaries in a given period range. This is in tension with the results for the $20001000<P_{orb}<$ 2000 range, as our model predicts more RG+BH binaries than Sun+BH binaries. This is because the empirical filter that removes long-period solutions probabilistically places an upper limit on the number of solutions. For sources that pass the $\texttt{ruwe}\ >1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ days}/P_{orb}$ filters, this limit is $\sim 4\%$ for $20001000<P_{orb}<$ 2000. Because of this, the number of RG+BHs in Table 3 is probably an overestimate.
In comparison, El-Badry et al. (2023a) estimated $\sim 40,000$ low mass star + BH binaries with $300100<P_{orb}<$ 300 days with the Galaxy, and an upper limit of 4,000 binaries for $1000400<P_{orb}<$ 1000 days, also assuming a logarithmically flat orbital period distribution. We suspect the difference is due to how El-Badry et al. (2023a) estimated $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$ to model the Gaia DR3 selection. Because of this, their effective search volume is smaller, leading to a higher occurrence rate than in this work.
Table 3: Number of black holes
$P_{orb}$ range (day) Prob. detect N($d<1$ kpc)
Sun-like star + BH
$400100<P_{orb}<$ 400 0.18 $5\pm 2$
$1000400<P_{orb}<$ 1000 0.44 $<2$
$20001000<P_{orb}<$ 2000 0.05 $<20$
Red giant + BH
$1000400<P_{orb}<$ 1000 0.50 $<2$
$20001000<P_{orb}<$ 2000 0.05 $20\pm 4$
Note. — Number of black holes within 1 kpc of the Sun. Upper limits and uncertainties are $1\sigma$. To convert these results to the number of black holes in the entire Milky Way, multiply the values in the N($d<1$ kpc) column by 1000.
6 Discussion
6.1 Discrepancy in eccentricity
Figure 11: Top panel: eccentricity distribution correction factor derived from the forward modeling. Bottom panel: NSS eccentricity distribution for Sun-like stars with orbital periods $1000100<P_{orb}<$ 1000 days, before and after correction. Note that we only plot orbits with $e>0.01$. The blue line shows a fit to the corrected distribution: a beta function with $\alpha=1.7$, $\beta=3.2$, rescaled with a minimum of 0.017 and width of 0.98. The ruwe  filter does not have a strong bias against eccentricity, so the raw and corrected distributions look very similar.
There is a serious discrepancy between the Gaia NSS astrometric orbit and forward modeled eccentricity distribution, with the forward model predicting significantly more eccentric binaries than are present in the Gaia astrometric orbits catalog. We hypothesize that the starting eccentricity distribution used in the forward model is incorrect and leads to this discrepancy.
The period, eccentricity and mass ratio distributions are generally well known for short period binary systems with $P_{\rm orb}\approx 0.1-50$ days (e.g., Moe & Di Stefano, 2015; Soszyński et al., 2016; Rowan et al., 2022). At longer orbital periods, the primary search techniques of radial velocities and astrometry are both biased against the detection and characterization of high eccentricity binaries. The $P_{\rm orb}-e$ relationship for $P_{\rm orb}\gtrsim 50$ day is generally modeled by extrapolation from a small number of moderate eccentricity systems, and is therefore poorly known and highly sensitive to the assumed completeness correction (Moe & Di Stefano, 2017). Because of these systematic biases, the intrinsic eccentricity distribution for wide binaries is highly uncertain.
Over time, the eccentricity distribution for sun-like stars with $10\lesssim P_{orb}\lesssim 1000$ days has been inferred to be more and more eccentric. Duquennoy & Mayor (1991) found a Gaussian eccentricity distribution with mean value $e\approx 0.3$. An updated study by Raghavan et al. (2010) instead found evidence for a flat eccentricity distribution. Moe & Di Stefano (2017) used the Raghavan et al. (2010) data to infer a power-law distribution in eccentricity $p(e)\propto e^{\eta}$, with $\eta=0.2\pm 0.3$ for $1000100<P_{orb}<$ 1000 days. Moe & Di Stefano (2017) then combined their result with other literature datasets and found for main sequence binaries with primary masses $0.8M_{\odot}<M_{1}<3M_{\odot}$ and orbital periods $6.00.5<\log P_{orb}<$ 6.0 days,
$\eta=0.6-\frac{0.7}{\log P_{orb}/\textrm{day}-0.5}$ (9)
(Equation 17 of Moe & Di Stefano 2017). This parametrization results in $\eta=0.13$ for $P_{orb}=100$ days and $\eta=0.32$ for $P_{orb}=1000$ days, and was the assumed eccentricity distribution used in the forward model presented in this work.
We use our forward model to determine a correction that can be applied to the observed NSS eccentricity distribution to convert it into the true underlying eccentricity distribution. We calculate the ratio of the normalized distribution of mock NSS astrometric orbits solution to the normalized distribution of the birth population as a function of eccentricity in four period bins. The reason for binning the results by period is that the eccentricity distribution is a function of period in the forward model (Equation 9). We then average the ratios across the period bins, and fit a quadratic through this average to determine a correction factor on the eccentricity distribution as a function of eccentricity (Figure 11). As expected, the astrometric cascade has a slight preference for filtering out high-eccentricity systems. This is due to the $\texttt{ruwe}>1.4$ filter (§4.1.1).
El-Badry et al. (2024) find a similar discrepancy between the Gaia NSS astrometric orbit and their forward modeled eccentricity distribution, with their forward model also predicting more eccentric binaries than are present in the Gaia astrometric orbits catalog. However, their discrepancy is not as significant as found in this work; they find their astrometric cascade produces a significantly stronger bias against the detection of eccentric orbits than we do. This bias is due to the eccentric orbits having larger astrometric uncertainties $\sigma_{a_{0}}$ because they are preferentially sampled near apastron; these solutions are then removed by the $a_{0}/\sigma_{a_{0}}$ filter. Although we do model the $a_{0}/\sigma_{a_{0}}$ filter in this work, our model for $\sigma_{a_{0}}$ is only a function of the number of visibility periods $n_{vis}$ and magnitude $G$ and does not account for the eccentricity or period of the orbit (§4.2). Thus, it does not capture the behavior observed in El-Badry et al. (2024), and the correction derived in Figure 11 thus only corrects the bias due to to the ruwe  filter.
We apply our correction to the eccentricity distribution in the NSS astrometric orbits catalog for Sun-like stars (which we roughly define as sources with $64<M_{G,0}<$ 6) and orbital periods $1000100<P_{orb}<$ 1000 (Figure 11) [^11] We note that that these two requirements do not actually change the shape of the eccentricity distribution as compared to the full NSS catalog, they just reduce the number of sources.. We find that the corrected distribution is very similar to the raw distribution (i.e., in our model, the detection probability for a binary is largely independent of eccentricity), and that the corrected Gaia NSS astrometric orbits catalog still contains significantly fewer high eccentricity binaries than predicted by Moe & Di Stefano (2017).
With Gaia DR4 per-epoch astrometry, these findings can be tested and validated. As a cross check, we also advocate for an independent method of measuring eccentricities that is less biased against high-eccentricity systems. Measuring the eccentricities of long-period eclipsing binaries from the Kepler survey (Kirk et al., 2016) is one way to accomplish this goal . Kepler continually monitored a single part of the sky over 4 years and is thus sensitive to eclipsing binaries with orbital periods of up to $\approx 2$ years.
6.2 Limitations of the approach
The filters described in §4.2 and 4.3 are empirical, and we do not attempt to derive the relations from first principles nor from the scanning law of the spacecraft. The empirical period filters attributed to the loss of orbital solutions assigned to acceleration solutions in the model cascade likely cannot be derived from first principles. In addition, it does not include potential correlations with other parameters (e.g., eccentricity) which might affect the probability of a source being given an acceleration solution. The empirical period filter is likely adequate for population studies, but less ideal when trying to infer properties of one single system. In addition, the empirical period filters do not map directly to a step in the model cascade, as the ruwe  and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters do. To do this would require forward modeling individual observations. For a more detailed approach that actually does simulate individual Gaia observations including the scanning law, see El-Badry et al. (2024) (further discussion is also presented in §6.5).
6.3 Spectroscopic solutions
Halbwachs et al. (2023) discuss in their §5.3 “Selection of the orbital solutions” that 2470 astrometric orbit solutions were recovered after being rejected by the post-processing by combining them with single-lined spectroscopic binary (SB1) solutions, and another 3500 solutions were discarded after post-processing due to validation described in Babusiaux et al. (2023). The processing description for the SB1 solutions was recently published (Gosset et al., 2024). These solutions represent about 3.5% of the total number of solutions in the NSS catalog, and so are likely negligible at the accuracy of the forward model presented in this work. However, it would be a limiting factor in achieving a precise understanding of selection effects and completeness; future work will incorporate the description of the SB1 processing. Ultimately, the epoch data of Gaia DR4 will mitigate this issue (§6.4).
6.4 Gaia DR4
As of late 2024, Gaia DR4 is projected to be released in mid-2026. DR4 will nearly double the time baseline of DR3 (66 months of data vs. 34 months). It is expected that the parallax and photocenter orbit semimajor axis uncertainties will be reduced by a factor of $\sqrt{2}$ due to having double the number of observations, and proper motion errors will be reduced by a factor of $\sqrt{8}$, with a factor of 2 in the doubled baseline, and another factor of $\sqrt{2}$ for the doubled number of observations.
It is beyond the scope of this work to extend this forward model framework to forecast for Gaia DR4. Several parts of this forward model, such as the empirical removal of solutions at certain orbital periods (§4.3), would require knowledge of the DR4 processing pipeline to forecast. An alternative way to create a conservative forecast of the selection of DR4 binaries is to model per-epoch photometry and assume that the processing cascade remains the same as that of DR3 (El-Badry et al., 2024). However, once DR4 is released, it would not be difficult to extend the model presented here.
Finally, we also note that DR4 will be the first Gaia data release to include individual epoch photometry and astrometry. Thus, the completeness and selection function can be evaluated in other manners, such as injection and recovery tests.
6.5 Comparison to El-Badry et al. (2024)
In order to characterize the selection function of the Gaia DR3 NSS astrometric orbits catalog, El-Badry et al. (2024) forward model individual epoch observations and validate it using a binary population synthesis model. Many of the underlying components of the population synthesis are similar or identical to that described in this paper; they also use galaxia and COSMIC to simulate the population of binary stars within 2 kpc. Minor differences include the extinction ($A_{G}=2.3E(B-V)$ in this work vs. $A_{G}=2.8E(B-V)$ in El-Badry et al. 2024), number of binaries at birth within 2 kpc (170 million in this work vs. 195 million in El-Badry et al. 2024), and the white dwarf initial-final mass relation (Kalirai et al. 2008 in this work vs. Weidemann 2000 in El-Badry et al. 2024).
To generate the individual observations, they use GOST to obtain the scan times and scan angles, and use Holl et al. (2023) for the noise model of the observations. They then fit these mock observations following the cascade described in Halbwachs et al. (2023). The advantage of modeling the observations at the epoch level is that each step described in the processing pipeline can be modeled. The main drawback is the computational expense. It requires approximately $10-20$ seconds to run the full astrometric cascade of El-Badry et al. (2024) for each source that receives an orbital solution, or about 5000 CPU hours to produce a full mock DR3 binary catalog. In contrast, it requires approximately 2 seconds to run 100,000 evaluations of the empirical forward model of this work at a fixed position on sky; most of that time comes from querying the 3-D dust map.
We also compare our results for the inclusion probability of Gaia BH1, BH2, and BH3 in the Gaia NSS astrometric orbits catalog. The results agree well for Gaia BH1 (33% in this work, and 39% in El-Badry et al. (2024)). For Gaia BH2 and BH3, El-Badry et al. (2024) find significantly higher inclusion probabilities (48% and 10%, as compared to the 4% and 0% in this work). One difference is that the values for El-Badry et al. are sky-averaged, while the values presented here are a function of (RA, Dec). For Gaia BH1, the detection probability would be 2×2$\times$2 × higher at higher ecliptic latitudes (Figure 9); for Gaia BH2, the detection probability does not vary across the sky (Figure 9). El-Badry et al. (2024) also calculate the change in inclusion probabilities as a function of distance. For Gaia BH1, the results are qualitatively similar, although our inclusion probability rises more steeply toward closer distances. For Gaia BH2, the results disagree, mainly because our inclusion probability does not exceed $\approx 4\%$. This may be due to strong covariances between $P_{orb}$, the mass of the dark companion, and detection probability, which is not captured in the probabilistic model that removes long orbital period solutions.
7 Conclusions
The Gaia DR3 NSS catalog has dramatically increased the number of well-characterized binary systems, and enabled the first mass measurement of black holes in detached binary systems. However, the selection function of this catalog is not well-characterized, and without knowledge of the selection function, discoveries of interesting systems are limited to individual curiosities, as their population statistics cannot be inferred.
To address this, we create a framework to forward model the selection function of the Gaia DR3 NSS astrometric orbits catalog. Using a binary population synthesis model, we validate this framework. We find the method produces reasonable agreement with the observed distribution of periods, parallaxes, photocenter sizes, ruwe, magnitudes, and parallax errors; the one parameter that is substantially discrepant is the eccentricity; we suspect the assumed birth population of eccentricities is the culprit.
As an example application of this framework, we use it the model to calculate the probability of Gaia BH1, BH2, and BH3 being included in the NSS astrometric orbits catalog. We find that Gaia BH1 has a relatively high chance of being included, but Gaia BH2 was a much luckier detection. Gaia BH3 was unlikely to be detected, due to its very long orbital period.
Ultimately, we look forward to Gaia DR4, where the time baseline will be doubled, and per-epoch astrometry will be released for all sources. This will enable proper characterization of the selection function and the discovery of many more interesting binary systems.
The code to calculate the probability of a binary being included in Gaia DR3 is available at https://github.com/caseylam/dr3_nss_ast_orbits_selection.
C.Y.L. acknowledges support from the Harrison and Carnegie Fellowships. K. E.-B. is supported by NSF grant AST-2307232. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has made use of NASA’s Astrophysics Data System.
Appendix A Parallax and photocenter orbit semimajor axis error model
Here we investigate the relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$, and $\sigma_{AL}/\sqrt{N_{vis}}$. There is a linear relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$ and $\sigma_{AL}/\sqrt{N_{vis}}$. Figure 12 shows this relationship, binned by the magnitude of the source and the number of visibility periods used. The binned value is indicated by the dot, the error bar is the error on the mean of that bin. Each dot represents a visibility period for that magnitude range; they are connected in order of decreasing number of visibility periods. The relationship gets noisy towards fainter magnitudes ($G\gtrsim 16$) and hence the line connecting the dot is no longer straight. The relationship for $138<G<$ 13 are all similar because $\sigma_{AL}$ is roughly constant in this magnitude range. However, it is unclear why going to fainter magnitudes, $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ are smaller for a given $\sigma_{AL}/\sqrt{N_{vis}}$. To investigate, we make the corresponding plot in $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ but using a random subset of the gaia source catalog. We find a similar relation persists, albeit the discrepancy at fixed $\sigma_{AL}/\sqrt{N_{vis}}$ is not as pronounced. We thus suspect that there is a magnitude dependence on $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ that is not necessarily entirely captured in $\sigma_{AL}$.
Figure 12: $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (top panel), $\sigma_{a_{0}}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (middle panel), and $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for a random subset of gaia_source sources (bottom panel). The sources are binned by magnitude and number of visibility periods used; for a given magnitude, the points are connected by a line indicated increasing number of visibility periods. The NSS catalogs exhibit a discontinuity between different magnitudes not seen in the gaia_source catalog.
Appendix B Cross-match to external catalogs
To test the hypothesis that the decrease in solutions with orbital periods $P_{orb}>550$ days is due to the model cascade and losing solutions to acceleration solutions, we cross-matched the Gaia DR3 NSS astrometric orbits and acceleration catalogs to two other binary star catalogs.
B.1 Cross-match to SB9
First, we cross-matched the SB9 catalog (Pourbaix et al., 2004) to the Gaia DR3 nss_acceleration_astro  and nss_two_body_orbit  catalogs to empirically determine the number of orbital solutions lost to the acceleration catalog. [^12] In §6 “Non-single stars” of Babusiaux et al. (2023), they perform the same exercise and find a “significant fraction” of astrometric acceleration solution could have been fit with an orbital solution, and that the acceleration values disagree from the known values; however, they do not specify how many or what fraction.
The top panel of Figure 13 shows the distribution of orbital periods from SB9, as well as the periods of astrometric acceleration and orbital solutions cross-matched to SB9. The middle panel is a histogram of the number of the cross-matched acceleration and orbital solutions in the SB9 catalog with $2000500<P_{orb}<$ 2000 days. The number of cross-matched acceleration solutions increases with increasing orbital period, while the number of cross-matched orbital solutions decreases; however, the number of cross-matched acceleration solutions increases slower than the number of cross-matched orbital solutions decreases, as their sum is not flat (middle panel of Figure 13).
However, the SB9 catalog is a highly heterogeneous catalog and the period distribution is not representative of an unbiased sample. To try and mitigate this effect, we also consider the number density of each type of solution; that is, the number of cross-matched orbital or cross-matched acceleration solutions at a given period, divided by the total number of SB9 orbital solutions at that period. The number density of cross-matched solutions is roughly flat (bottom panel of Figure 13).
We re-emphasize that the SB9 catalog is a heterogeneous and biased sample, so detailed conclusions cannot be drawn from this test. However, this test suggests that our hypothesis that the decrease in solutions with $1000550<P_{orb}<$ 1000 days is due to the model cascade is plausible.
B.2 Cross-match to WDS-ORB6
Next, we perform the same cross-matching exercise instead using the Sixth Catalog of Orbits of Visual Binary Stars (Orb6, an updated version of Hartkopf et al. (2001)). We used the cross-match of Orb6 to Gaia EDR3 from Chulkov & Malkov (2022), then matched their Gaia EDR3 source IDs to the NSS catalogs. Orb6 binaries are graded based on the reliability of the solution as well as the detection method. We only include binaries that have grades 1 (“definitive”) or 2 (“good”), or are detected astrometrically or interferometrically.
The same notes of caution for the SB9 catalog also applies here, although the results are even more tentative here, as the Orb6 orbits are substantially less reliable than the SB9 orbits, even after applying our quality cuts based on cross-match of orbital solutions to available Gaia astrometric orbits.
Figure 13: Top panel: period distribution of SB9 orbits (blue), SB9 orbits cross-matched to the NSS astrometric orbit solutions (green), and SB9 orbits cross-matched to the NSS astrometric acceleration solutions (orange). Middle panel: Number of NSS astrometric acceleration solutions (blue dot), NSS astrometric orbit solutions (orange dash), and their sum (green solid). Bottom panel: Same as middle panel, but normalized by the number of total SB9 solutions.
Figure 14: Same as Figure 13, but for the Orb6 catalog instead of the SB9 catalog.

### v2

@@
 In June 2022, the Gaia mission released a catalog of astrometric orbital solutions for 168,065 binary systems, by far the largest such catalog to date.
-Unlike previous binary stars catalogs, which were heterogeneous collections of orbits from different surveys and instruments, these orbits were derived using Gaia data alone.
-Despite this homogeneity, the selection function is difficult to characterize because of choices made in the construction of the catalog.
+The catalog’s selection function is difficult to characterize because of choices made in its construction.
 Understanding the catalog’s selection function is required to model and interpret its contents.
 We use a combination of analytic and empirical prescriptions to construct a function that computes the probability that a binary with a given set of properties would have been published in the Gaia Data Release 3 astrometric orbit catalog.
-We also construct a binary population synthesis model based on Moe & Di Stefano (2017) to validate our characterization of the selection function, finding good agreement with the actual Gaia NSS catalog, with the exception of the orbital eccentricity distribution.
+This is a complementary approach to the more accurate but significantly more computationally expensive approach of El-Badry et al.
+(2024).
+We also construct a binary population synthesis model to validate our characterization of the selection function, finding good agreement with the actual Gaia NSS catalog, with the exception of the orbital eccentricity distribution.
 The NSS catalog suggests high-eccentricity orbits are relatively uncommon at intermediate periods $100\lesssim P_{orb}\lesssim 1000$ days.
-As an example application of the selection function, we estimate the Gaia DR3 detection probabilities of the star + BH binaries Gaia BH1, BH2, and BH3.
-We also estimate the population of Sun-like star + BH binaries in the Galaxy to be $\sim 5000$ for $400100<P_{orb}<$ 400 day, $\lesssim 2,000$ for $1000400<P_{orb}<$ 1000 day, and $\lesssim 20,000$ for $20001000<P_{orb}<$ 2000 days.
-facilities: Gaiasoftware: isochrones (Morton, 2015), galaxia (Sharma et al., 2011), astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2019), healpy (Zonca et al., 2019), mwdust (Bovy et al., 2019), COSMIC (Breivik et al., 2020)
+As an example application of the selection function, we estimate the Gaia DR3 detection probabilities of the star + BH binaries Gaia BH1 and BH2, and find them to be 0.38 and 0.27, respectively.
+Compared to the values obtained by detailed modeling in El-Badry et al.
+(2024), the probabilities are identical for BH1, and within a factor of 2 for BH2.
+We also estimate the population of Sun-like star + BH binaries in the Galaxy to be $\sim 3000$ for $400100<P_{orb}<$ 400 day, $<800$ for $1000400<P_{orb}<$ 1000 day, and $<12,000$ for $15001000<P_{orb}<$ 1500 day.
+facilities: Gaiasoftware: isochrones (Morton, 2015), galaxia (Sharma et al., 2011), astropy (Astropy Collaboration et al., 2013, 2018, 2022), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2019), healpy (Zonca et al., 2019), mwdust (Bovy et al., 2019), COSMIC (Breivik et al., 2020)
 1 Introduction
 Gaia is an astrometric space mission surveying over 2 billion stars in the Milky Way (Gaia Collaboration et al., 2016a).
-Its main mission operated from July 2014 through December 2019, and its extended mission is planned to operate through January 2025.
+Its main mission operated from July 2014 through December 2019, and its extended mission operated through January 2025.
 The data products and catalogs generated from the photometry, astrometry, and spectroscopy from Gaia are shared publicly in data releases that occur every several years.
@@
 We focus on the solutions that have an astrometric orbit (as compared to those that only have a spectroscopic orbit) as our primary goal is to understand the selection function of stellar-mass black hole + star binaries, which have thus far only been found in the astrometric orbits catalog.
-Our forward modeling approach uses analytic and simple prescriptions to approximate the steps in the construction of the catalog.
-This is a complementary approach to that of El-Badry et al.
-(2024), who forward model individual observations at the expense of computational cost; more discussion on the tradeoffs between the two approaches is presented in §6.5.
+However, we emphasize that the selection function is applicable to any type of binary system, e.g., star + star, or star + planet.
+The selection function is defined in terms of observables (e.g., the size of the photocenter’s orbit; Equation 1).
+Although different companion luminosities and masses will result in different observables and thus different detection probabilities, Gaia is agnostic to whether the secondary component of an unresolved binary is a star, planet, or black hole.
+The purpose of the selection function is to quantify how different those detection probabilities are.
+Our forward modeling approach combines analytic models and results from the individual-epoch forward modeling of El-Badry et al.
+(2024) to produce fast, empirical functions to calculate the probability of a source being included in the NSS astrometric orbits catalog.
 The rest of this paper is organized as follows.
 In §2 we review the processing pipeline of the Gaia NSS astrometric orbit catalog and identify the most important steps to model.
-In §3 we describe the binary population synthesis model.
+In §3 we describe the binary population synthesis model that we compare to the NSS catalog.
 In §4 we forward model the NSS astrometric processing cascade, and the results are presented in §5.
-Applications, shortcomings, possible extensions of the forward model, and comparison to other work are discussed in §6.
+Applications, strengths and shortcomings, and possible extensions of the forward model are discussed in §6.
 We provide a final summary and conclusions in §7.
@@
 This resulted in a series of many filtering steps to remove potentially spurious solutions, at the expense of many accurate solutions.
-Our goal is to understand the selection effects of this filtering process.
+Our goal is to understand the selection effects of this filtering process, as well as determine which filters are most important to model.
 The astrometric NSS processing described in Halbwachs et al.
-(2023) is summarized below, and the portions most relevant to the forward modeling are highlighted; see also Figure 1 of Halbwachs et al.
+(2023) is summarized below, and the portions most relevant to the forward modeling are highlighted in Figure 1; see also Figure 1 of Halbwachs et al.
 (2023) for a summary.
+Figure 1: Summary of the NSS astrometric orbit cascade, illustrating the criteria that must be satisfied for a source to be included in the NSS astrometric orbit catalog.
 2.1 Astrometric orbit solutions
@@
 This paper is only concerned with modeling the subset of solutions that have astrometric orbits, i.e.
-Orbital and AstroSpectroSB1 solutions from the the nss_two_body_orbit table.
+Orbital and AstroSpectroSB1 solutions from the nss_two_body_orbit table.
 We collectively refer to these as the NSS astrometric orbit solutions.
+Unless otherwise stated, we will also just use the shorthand “NSS solutions” or “NSS catalog” to describe this dataset, since we are only concerned with astrometric models in this paper.
 2.2 Selection of stars to be processed
-An initial filtering was performed on the entire Gaia DR3 gaia_source  catalog to define the input catalog, a subset of sources that were good NSS candidates.
+In the astrometric NSS processing, an initial filtering was performed on the entire Gaia DR3 gaia_source  catalog to define the input catalog, a subset of sources that were good NSS candidates.
 Starting from $1.5\times 10^{9}$ sources with single-star astrometric solutions, several filtering steps were implemented to produce the input catalog of $4.1\times 10^{6}$ stars.
@@
 As described in Halbwachs et al.
-(2023) §2.1, the filters are as follows:
-[…]
-  • $\texttt{ruwe}>1.4$ (square root of the reduced $\chi^{2}$, with an empirical renormalization based on magnitude and color of source),
-[…]
-[…]
-[…]
-  • $C^{*}<1.645\sigma_{C^{*}}$ (color excess (ratio between the total flux in BP and RP, and the G-band flux (BP + RP)/G) after correcting for passbands, possibly from a contaminating neighbor).
-We only model the ruwe  filter.
-The ipd_frac_multi_peak, ipd_gof_harmonic_amplitude, and $C^{*}$ filters are intended to remove marginally resolved wide binaries.
-Such binaries tend to have orbital periods $P_{orb}\gtrsim$ decades, which are much too long to receive an astrometric orbital solution; they have $\texttt{ruwe}>1.4$ because their marginally resolved nature leads to spurious epoch astrometry.
-The filters on apparent magnitude and visibility periods are accounted for in the $\texttt{ruwe}>1.4$ filter.
+(2023) §2.1, the filters are on the source’s apparent magnitude, goodness of fit of a single star solution, number of visibility periods (i.e., independent observations), [^1] Using the number of visibility periods provides a more reliable estimate for position and proper motion uncertainties than the total number of good along-scan observations (Lindegren et al., 2018).
+image parameter determination metrics, and color excess.
+The filters on the image parameter determination metrics and color excess are designed to remove marginally resolved wide binaries.
+Gaia DR3 has an effective angular resolution of around 0.7 arcsec [^2] Although Gaia DR3 has an angular resolution of 0.18 arcsec (Lindegren et al., 2021), it is limited spatially by crowding, making its effective angular resolution 0.7 arcsec.; at smaller separations the completeness of close source pairs drops (Gaia Collaboration et al., 2021a).
+In our forward model, we will only model the filter on the goodness of fit of a single star solution ($\texttt{ruwe}>1.4$, §4.2).
+The filters on apparent magnitude and visibility periods are accounted for in the ruwe  filter.
 As ruwe  is a measurement of the significance of deviations from a single-star model, sources that are faint or have too few visibility periods will have astrometric uncertainties too large to have a significant deviation.
-2.3 Model cascade
-Subsequently, the $4.1\times 10^{6}$ stars from the input catalog were fit with a variety of astrometric NSS models. The models were tried in the following order:
+In addition, our forward model does not explicitly model the image parameter determination metrics or color excess, because the wide binaries removed by these sources will be removed by the $\texttt{ruwe}>1.4$ filter.
+Using a binary population synthesis model (§3), we find this angular resolution corresponds to binaries within 2 kpc having orbital periods several centuries to millennia.
+These orbital periods are much too long to receive an astrometric orbital solution; they have $\texttt{ruwe}>1.4$ because their marginally resolved nature leads to spurious epoch astrometry.
+See also Tokovinin (2023) and Appendix B of El-Badry et al.
+(2024).
+2.3 Model cascade and main processing
+Subsequently in the astrometric NSS processing pipeline, the $4.1\times 10^{6}$ stars from the input catalog were fit with a variety of astrometric NSS models.
+The models were tried in the following order [^3] We do not model the “variability-induced movers” branch in the processing cascade; this describes sources with astrometric motion in unresolved binaries where one source is photometrically variable.
+We do not consider variable sources in our catalog so this is not relevant.:
   • variable acceleration (Acceleration9 solution, nss_acceleration_astro table; 9 parameters),
   • constant acceleration (Acceleration7 solution, nss_acceleration_astro table; 7 parameters),
-  • orbital (Orbital solution, nss_two_body_orbit table; 12 parameters11110 parameters were used for a circular or psuedo-circular orbit (see §5.2 “Pseudo-circular and circular orbital solutions” of Halbwachs et al.
+  • orbital (Orbital solution, nss_two_body_orbit table; 12 parameters44410 parameters were used for a circular or psuedo-circular orbit (see §5.2 “Pseudo-circular and circular orbital solutions” of Halbwachs et al.
 (2023)).
@@
 (2023)).).
+In the NSS astrometric processing pipeline, there are two steps: main processing and post-processing.
+If a source passes a given model’s main processing criteria, it moves into the post-processing branch of that model; if a source does not pass a given model’s main processing criteria, the next model in the cascade is tried.
+If a source passes both a given model’s main processing and post-processing criteria, it is accepted as a solution and is included in that model’s catalog.
+If a source does not pass the post-processing criteria, it is removed and given a single star solution.
+For each model, the criteria for main processing acceptance were based on the the significance of the model and the significance of the parallax measurement.
+The specific criteria are summarized in Table 1 of Halbwachs et al.
+(2023).
 If an acceptable solution was found, subsequent models were not tried.
@@
 Thus, if an acceptable acceleration model was found, an orbital model was not tried, even if an orbital solution would have led to a better fit.
-We discuss the implications of this choice in §4.3.
+In our forward model, we model the variable acceleration (referred to as 9-parameter acceleration), constant acceleration (referred to as the 7-parameter acceleration), and the orbital components of the astrometric cascade.
+For each of these three models, we model the main processing filters on the significance of the model and the parallax measurement.
+Throughout this paper we discuss acceptance of acceleration and orbital models.
+When we say an acceleration model is accepted or rejected, we are specifically referring to the main processing step unless otherwise stated.
+This is because only the main processing branch of the acceleration models is relevant to the orbital solutions.
+If an acceleration model is rejected from the main processing, it is a candidate for an orbital solution.
+If an acceleration model is accepted in the main processing, it will not be a candidate for an orbital solution regardless of whether it is accepted or not in the subsequent acceleration post-processing.
+In contrast, when we say an orbital model is accepted or rejected, we are specifically referring to acceptance in both the main and post-processing steps.
+This is because we are interested in the final astrometric orbits catalog, which requires passing the post-processing criteria.
 2.4 Orbital solution model acceptance
-For sources where an orbital model fit was attempted, there were several conditions determining whether the solution was accepted and included in the final orbital solution catalog.
-A stringent filter was applied on the significance of the parallax solution ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$) where $\varpi$ is the parallax, $\sigma_{\varpi}$ is the parallax error, and $P_{orb}$ is the orbital period.
-This filter was designed to remove false positives with spurious astrometric orbits and orbital periods related to the Gaia scanning law (Holl et al., 2023), but came at the expense of many valid solutions.
-There were also several other less stringent filters.
-A filter was applied on the significance of the astrometric orbit $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ where $a_{0}$ and $\sigma_{a_{0}}$ are the photocenter orbit semimajor axis and its uncertainty.
-Another filter was applied on the error of the orbital eccentricity $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$.
-Finally, a filter was also applied on the goodness of fit of the binary solution $F_{2}=\sqrt{9\nu/2}[(\chi^{2}/\nu)^{1/3}+(2/9\nu)-1]<25$, where $\chi^{2}$ is the sum of the squared normalized residuals and $\nu$ is the number of degrees of freedom.
-We only model the filters on $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$.
-The filter on $\sigma_{e}$ only removes a small percentage of solutions.
-As described in §5.3 “Selection of the orbital solutions” of Halbwachs et al.
-(2023), approximately 180,000 orbital solutions remained after applying the filters $F_{2}<25$, $a_{0}/\sigma_{a_{0}}>5$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$; subsequently applying the filters on $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$ and $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{day}}$ only removed an additional 7% of solutions.
-Thus, the contribution of the $\sigma_{e}$ filter cannot be more than 10% of the final NSS astrometric orbit catalog [^2] Using the modeling of El-Badry et al.
-(2024), we find that the $\sigma_{e}$ filter only removes 2% of sources, which is 1/3 of the 7% removed by the $a_{0}/\sigma_{a_{0}}$ filter.; this is negligible given the other uncertainties in the forward model.
-Visual inspection of Figure 3c of Halbwachs et al.
-(2023) also suggests the $\sigma_{e}$ filter has minimal impact on the final constructed catalog.
-We also do not model the filtering on $F_{2}$ because we assume that good orbital solutions will have good fits.
+For sources in the NSS astrometric processing pipeline where an orbital model fit was attempted in the main processing, there are three additional conditions determining whether the solution was accepted in the post-processing and included in the final orbital solution catalog: a criterion on the goodness-of-fit $(F_{2})$, a second more stringent filter on the significance of the model, and a filter on the error of the eccentricity measurement $\sigma_{e}$ (Halbwachs et al., 2023).
+In our forward model, we do not model the filter on $\sigma_{e}$, as it removes no more than a few percent of solutions and does not bias the inferred distributions of orbital parameters.
+Detailed justification is provided in Appendix A.
+Visual inspection of Figure 3 of Halbwachs et al.
+(2023) also suggests the filter on $\sigma_{e}$ has minimal impact on the final constructed catalog.
+In our forward model, we also do not model the filtering on $F_{2}$ because we assume that good orbital solutions will have good fits.
 $F_{2}$ is expected to follow a normal distribution with mean 0 and standard deviation 1; even if the uncertainties are poorly characterized, $F_{2}<25$ is still a 25-$\sigma$ threshold.
 It may also allow false positives to pass through, but is unlikely to remove real solutions that have passed the other filtering steps (with the exception of binaries in hierarchical triple systems).
+We provide further justification of this choice in Appendix B.
 2.5 Final NSS astrometric orbit catalog
-The black outline in Figure 7 shows the distribution of orbital period, parallax, eccentricity, photocenter orbit semimajor axis, renormalized unit weight error, apparent magnitude, and parallax over parallax error for the final set of 168,065 NSS astrometric orbit solutions.
+Figure 2: 1-D and 2-D PDFs in the period $P_{orb}$, parallax $\varpi$, eccentricity $e$, photocenter orbit semimajor axis $a_{0}$, ruwe, apparent magnitude $G$, and parallax over parallax error $\varpi/\sigma_{\varpi}$ of the actual NSS astrometric orbits catalog (black) as compared to forward model (red, described in §35).
+In general, the forward modeling does a good to reasonable job in matching the actual NSS distributions including the correlations, with the exception of the eccentricity distribution.
+The black outline in Figure 2 shows the distribution of orbital period, parallax, eccentricity, photocenter orbit semimajor axis, renormalized unit weight error, apparent magnitude, and parallax over parallax error for the final set of 168,065 Gaia DR3 NSS astrometric orbit solutions.
 Some of the correlations are clearly due to the filtering procedure (e.g.
 orbital period vs.
-parallax over parallax error.)
+parallax over parallax error).
 3 Binary population synthesis
@@
 Note that only binary stars are modeled; single stars are not included, nor are higher-order multiple systems.
+We emphasize that the population synthesis is used primarily to validate the output of the selection function, and that it is not used in determining the selection function itself.
+Once the selection function has been determined, different population synthesis models can be explored to determine whether they produce different predicted numbers of detectable black holes (e.g., Nagarajan et al., 2025), although such a study is out of the scope of the work presented here.
 3.1 Galactic model (positions, kinematics, ages, metallicities, extinction)
@@
 Hence, we combine galaxia positions, metallicities, and ages with a binary star population based on Moe & Di Stefano (2017), which is a joint probability distribution in primary mass $M_{1}$, mass ratio $q=M_{2}/M_{1}$, orbital period, and eccentricity.
-Moe & Di Stefano (2017) infer the distribution bover $0.8<M_{1}<40M_{\odot}$, $10.1<q<$ 1, $80.15<\log P_{orb}/\textrm{day}<$ 8, and $10<e<$ 1.
+Moe & Di Stefano (2017) infer the distribution over $0.8<M_{1}<40M_{\odot}$, $10.1<q<$ 1, $80.15<\log P_{orb}/\textrm{day}<$ 8, and $10<e<$ 1.
 For reference, about 60% of Sun-like stars are single, while this decreases to nearly 0% for OB-type stars.
@@
 This is because the binary statistics in Moe & Di Stefano (2017) only extend down to $0.8M_{\odot}$; to extrapolate the multiplicity statistics down to $0.08M_{\odot}$, COSMIC assumes the binary fraction decreases linearly from $\approx 40$% at $0.8M_{\odot}$ down to 0 at $0.08M_{\odot}$.
-This choice of implementation is somewhat arbitrary (given that M-dwarfs appear to have a binary frequency $\sim 20\%$; e.g., Laithwaite & Warren 2020), but ultimately these sources are too faint to be included in the astrometric orbits catalog, so their assumed binary fraction does not matter.
-Assuming that the majority of binary systems are unresolved, one galaxia star corresponds to one COSMIC stellar system.
+This choice of implementation is somewhat arbitrary (given that M-dwarfs appear to have a binary frequency $\sim 20\%$; e.g., Laithwaite & Warren 2020), but ultimately most of these sources are too faint to be included in the astrometric orbits catalog, so the model is not sensitive to the particular binary fraction assumed.
+Based on an estimate using the extinction-corrected absolute magnitudes of the NSS sources, we find that main sequence M-dwarf binaries comprise $\lesssim 5\%$ of the NSS catalog.
+Similarly, we do not simulate substellar companions, as they are a negligible part of the NSS catalog; Gaia Collaboration et al.
+(2023b) find that only 1% of NSS companions are brown dwarfs or exoplanets.
+Assuming that the majority of binary systems are unresolved [^8] Using the criterion for resolvability given by Equation (B1) in El-Badry et al.
+(2024) and assuming the binary separation is equal to the semimajor axis, we find that 80% of binary star systems will be unresolved., one galaxia star corresponds to one COSMIC stellar system.
 This corresponds to 150 million binary systems within 2 kpc at the present day.
@@
 (2008).
-The remaining white dwarf + star binaries are assumed to either have orbital periods too short ($\lesssim 10$ days) or too long ($\gtrsim 10^{4}$ days) that they could not be detected astrometrically by Gaia in DR3.
+The remaining white dwarf + star binaries are assumed to either have orbital periods too short ($\lesssim 10$ days) or too long ($\gtrsim 10^{4}$ days) to be detected astrometrically by Gaia in DR3.
 We also remove binaries where either star is currently or has previously overflowed its Roche lobe, under the assumption that the periods will be too short to detect astrometrically in Gaia DR3 due to orbital shrinking.
@@
 Thus, $q_{l}=q$ if component 1 is both brighter and more massive, otherwise $q_{l}=1/q$.
-Note that the photocenter orbit semimajor axis is maximized for dark companions (i.e., $l=0$) and large mass ratios (i.e., $q_{l}\gg 1$), and is zero for equal-mass companions of the same luminosity ($l=1$, $q_{l}=1$).
-Finally, we note that all these luminosity-dependent quantities are specifically in the Gaia G-band.
+The photocenter orbit semimajor axis is maximized for dark companions (i.e., $l=0$) and large mass ratios (i.e., $q_{l}\gg 1$), and is zero for equal-mass companions of the same luminosity ($l=1$, $q_{l}=1$).
+We note that all these luminosity-dependent quantities are specifically in the Gaia G-band.
+3.3 Gaia-specific observational quantities
+Finally, the population synthesis model also requires information about Gaia’s observational properties: the number of observations and the measurement error.
+To obtain an estimate for the number of visibility periods $N_{vis}$ in the population synthesis, we use the Gaia Observation Forecast Tool (GOST) [^10] https://gaia.esac.esa.int/gost/, which predicts when any given part of the sky crosses Gaia’s focal plane.
+To calculate $N_{vis}$, we evaluate the number of predicted scans at 49152 evenly spaced points on the sky, discard a random 10% of the predictions to mimic losses due to dead time [^11] The GOST page suggests 20% would be more appropriate, but we find 10% gives good agreement with the Gaia DR3 catalog., then calculate how many of the remaining predicted observations are separated by at least 4 days.
+We adopt the along-scan per-CCD centroiding error $\sigma_{AL}(G)$ from Figure 3 of Holl et al.
+(2023).
+$\sigma_{AL}(G)$ is a function of the source magnitude $G$.
+As described in Lindegren (2018), for $12\lesssim G\lesssim 17$ the formal centroiding precision is limited by photon noise (leading to Poisson errors), for $G\gtrsim 17$ background noise dominates and the error rises more quickly than Poisson, and for $G\lesssim 12$ the CCD gates limit the number of photons and the centroiding precision is independent of magnitude.
+The $\sigma_{AL}(G)$ we adopt from Holl et al.
+(2023) is empirically determined from Gaia DR3 data, and is significantly larger ($>0.3$ dex) than the formal error for bright ($G\gtrsim 12)$ sources.
 4 Empirical modeling of Gaia NSS model cascade
 With a binary population model in hand, the next step is to reproduce the modeling cascade described in §2.
-In this section, we describe how we estimate and implement the filters $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$.
-In addition, we also need a way to account for the fact that 1) Gaia DR3 is based on only 34 months of data, 2) orbital solutions are “lost” higher up in the model cascade when they are fit with an acceleration solution, and 3) orbital periods of 1 year are very difficult to distinguish from parallax.
-We describe the empirical filters we apply to account for these effects.
-4.1 Empirical ruwe  model
+In this section, we describe how we estimate and implement the filter that selects astrometric NSS candidates ($\texttt{ruwe}>1.4$), the effect of the acceleration models in the cascade, and the filters that select the accepted orbital solutions ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ and $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$).
+Our general framework is to estimate the quantities of interest analytically, then use the simulations of El-Badry et al.
+(2024) to calibrate these estimates as a function of intrinsic orbital and binary parameters (e.g., period, ecccentricity, photocenter orbit size).
+4.1 El-Badry et al. (2024) simulations
+El-Badry et al.
+(2024) generate a synthetic population of Galactic binary stars, then forward model individual 1-D astrometric measurements of these sources according to Gaia’s scanning law and noise properties.
+They then model these 1-D observations according to the NSS astrometric cascade, fitting the simulated epoch data with the various NSS astrometric models and applying the various filtering criteria.
+El-Badry et al.
+(2024) find good agreement between the properties of their forward modeled catalog and the actual Gaia DR3 NSS astrometric orbits catalog.
+Many of the underlying components of the population synthesis in El-Badry et al.
+(2024) are similar or identical to those described in this paper; they also use galaxia and COSMIC to simulate the population of binary stars within 2 kpc.
+Minor differences include the extinction ($A_{G}=2.3E(B-V)$ in this work vs.
+$A_{G}=2.8E(B-V)$ in El-Badry et al.
+2024), number of binaries at birth within 2 kpc (170 million in this work vs.
+195 million in El-Badry et al.
+2024), and the white dwarf initial-final mass relation (Kalirai et al.
+2008 in this work vs.
+Weidemann 2000 in El-Badry et al.
+2024).
+However, we emphasize that the results we derive from El-Badry et al.
+(2024) in the rest of this section are independent of the details of their population synthesis.
+This is because we are considering the effects of the cascade relative to the initial catalog.
+As long as all relevant regions of parameter space are covered, then the results are independent of the specific assumptions used to generate the binary population.
+4.2 ruwe  model
 The renormalized unit weight error ruwe  is used to quantify the goodness-of-fit of a single star model.
@@
 If $\delta\theta$ is the astrometric excess noise due to fitting an unresolved binary with a single star model, then
-$\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1},$ (3)
-where $\sigma_{AL}(G)$ is the along-scan centroiding error, which is a function of the source magnitude $G$.
-Here, as well as in the remainder of this paper, we adopt the per-CCD centroiding error from Figure 3 of Holl et al.
-(2023).
-As described in Lindegren (2018), for $12\lesssim G\lesssim 17$ the formal centroiding precision is limited by photon noise (leading to Poisson errors), for $G\gtrsim 17$ background noise dominates and the error rises more quickly than Poisson, and for $G\lesssim 12$ the CCD gates limit the number of photons and the centroiding precision is independent of magnitude.
-The $\sigma_{AL}(G)$ we adopt from Holl et al.
-(2023) is empirically determined from Gaia DR3 data, and is significantly larger ($>0.3$ dex) than the formal error for bright ($G\gtrsim 12)$ sources.
+$\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1}.$ (3)
 […]
@@
 (2020) derived an analytic relationship (their Equation 17) between the astrometric excess noise $\delta\theta$, the source magnitude $G$, photocenter orbit semimajor axis $a_{0}$, inclination $i$, and eccentricity $e$
-$\delta\theta=a_{0}\sqrt{1-\frac{\sin^{2}i}{2}-\frac{3+\sin^{2}i(\cos^{2}\phi_{\nu}-2)}{4}e^{2}},$ (4)
-where $\phi_{\nu}$ is the azimuthal viewing angle of the binary.
+$\delta\theta=a_{0}\sqrt{1-\frac{\sin^{2}i}{2}-\frac{3+\sin^{2}i(\cos^{2}\omega-2)}{4}e^{2}},$ (4)
+where $\omega$ is the argument of periapse.
+[^12] In Penoyre et al.
+(2020), $\omega$ is denoted $\phi_{\nu}$ and is called the azimuthal viewing angle of the binary.
+We have chosen to use the more standard notation of the Campbell elements.
 Note that the derivation of Equation 4 assumes that there are many observations sampling the full orbital period of the binary during the survey, such that the observations can be time-averaged over the orbit.
-For an isotropic distribution of $\phi_{\nu}$, the average value of $\cos^{2}\phi_{\nu}$ is 0.5. Fixing $\cos^{2}\phi_{\nu}$ to 0.5, the astrometric excess noise can be re-written
+For an isotropic distribution of $\omega$, the average value of $\cos^{2}$$\omega$ is 0.5. Fixing $\cos^{2}$ $\omega$ to 0.5, the astrometric excess noise can be re-written
 $\delta\theta=a_{0}\sqrt{\frac{1}{2}\left(1+\cos^{2}i\right)\left(1-\frac{3e^{2}}{4}\right)}.$ (5)
@@
 There are still some minor trends visible in $|\cos i|$, but Gaia Collaboration et al.
-(2023b) showed that the sensitivity of circular orbits deviates slightly from the the expected $1+\cos^{2}i$, so this is not surprising.
+(2023b) showed that the sensitivity of circular orbits deviates slightly from the expected $1+\cos^{2}i$, so this is not surprising.
 Note that there are very few red points in the bottom panels because there are very few high-eccentricity orbits.
@@
 The predicted astrometric excess noise in Equation 5 and ruwe  in Equation 6 will generally be overestimates of the astrometric excess noise and ruwe  reported by Gaia.
-The derivation of Equation 5 assumes that the binary signal has not been erroneously absorbed into other astrometric terms (i.e., position, parallax, proper motion, or satellite pointing noise).
+The derivation of Equation 5 assumes that the binary signal has not been erroneously attributed to other astrometric terms (i.e., position, parallax, proper motion, or satellite pointing noise).
 Penoyre et al.
-(2020) predicted that this erroneous absorption would happen frequently, often lowering the reported astrometric excess noise and ruwe.
-With the NSS catalog, we confirm that the predicted astrometric excess noise $\delta\theta$ tends to be smaller than the actual $\delta\theta$ in the catalog by $\approx 60-70\%$ (Figure 3).
-We calculate an empirical correction to rescale the analytic ruwe  prediction to match the actual ruwe  using the NSS catalog.
-The number of visibility periods used in the solution ($n_{vis}$) affects the size of the erroneous absorption.
-For smaller $n_{vis}$, the average size of the erroneous absorption is larger, and there is more scatter about this average; for larger $n_{vis}$, the average size of the erroneous absorption is smaller, and there is less scatter about this average (Figure 3).
-To calculate the correction factor, we split the NSS astrometric orbit solutions into three bins based on the number of visibility periods used: $n_{vis}<23$, $2523\leq n_{vis}<$ 25, and $n_{vis}\geq 25$.
+(2020) predicted that this erroneous attribution would happen frequently, often lowering the reported astrometric excess noise and ruwe.
+With the NSS catalog, we confirm that the predicted astrometric excess noise $\delta\theta$ tends to be larger than the actual $\delta\theta$ in the catalog by $\approx 60-70\%$.
+We calculate an empirical correction to rescale the analytic ruwe  prediction to match the ruwe  measured by fitting a single-star model to the simulated observations in the catalog of El-Badry et al.
+(2024).
+We use a catalog that includes a random subset of the birth population of binaries for which the orbital parameters are known and ruwe is calculated, but before the $\texttt{ruwe}>1.4$ filter has been applied.
+For each simulated source in this catalog, we calculate $\delta\theta_{analytic}$ (the expected astrometric excess noise using Equation 5) and $\delta\theta_{fit}$ (the astrometric excess noise inferred from the ruwe  measured in the fitting procedure; Equation 3).
+The ratio $\delta\theta_{fit}/\delta\theta_{analytic}$ is the rescaling factor required to correct the analytic prediction of $\delta\theta$ to account for the erroneous attribution that occurs in the fitting process.
+Within each bin, we only consider sources with $\sigma_{AL}<a_{0}$ (as otherwise their orbits are not detectable), and sources with $\texttt{ruwe}>1.01$ (as otherwise the predicted value of $\delta\theta$ will be imaginary).
+In order to account for the dependence of this correction factor on the orbital period and number of observations, we split the simulated catalog solutions from El-Badry et al.
+(2024) into orbital period bins of 25 days, spanning 0 to 1500 days.
+We further subdivide each period bin based on the number of visibility periods used: $N_{vis}<25$, and $N_{vis}\geq 25$.
 These bins were chosen since $N_{vis}$ is a bimodal distribution for the NSS astrometric orbits catalog, with peaks at $N_{vis}=18$ and 27.
-In each of these bins, a Gaussian is fit to the distribution of the ratio of the actual $\delta\theta$ to the analytically predicted $\delta\theta$ from Equation 5.
-The mean $\mu$ and standard deviation $\sigma$ for each of the bins are $\mu=0.574$, $\sigma=0.133$ for $n_{vis}<23$; $\mu=0.606$, $\sigma=0.116$ for $2523\leq n_{vis}<$ 25, and $\mu=0.643$, $\sigma=0.099$ for $n_{vis}\geq 25$.
-To apply this correction, a sample is drawn from the appropriate Gaussian and used to rescale the analytic estimate $\delta\theta$, and in turn ruwe, since $\texttt{ruwe}=\sqrt{1+(\delta\theta/\sigma_{AL})^{2}}$ (Equations 4 and 5).
-We note that we chose not to directly calculate the correction using the ruwe  reported in the NSS astrometric catalog, because the hard limit $\texttt{ruwe}>1.4$ makes the distribution non-Gaussian and would be more difficult to model than using $\delta\theta$.
-Figure 3: Comparison of the astrometric excess noise $\delta\theta$ for sources in the NSS astrometric orbit catalog, vs.
-the analytic predicted $\delta\theta$ of those sources calculated using the photocenter orbit semimajor axis, eccentricity, and inclination from the NSS catalog and Equation 5.
-We perform the comparison by the number of visibility periods used $n_{vis}$ in the solution.
-The top panels show scatter plots of the actual $\delta\theta$ vs.
-analytic predicted $\delta\theta$, with the 1:1 relation is denoted by the orange line.
-The bottom panels show the ratio of the actual to analytic predicted $\delta\theta$, with the best-fit Gaussian shown in the black line.
-The predicted astrometric excess noise is systematically higher than the actual value reported in the NSS catalog.
-The actual values reported are typically only $\approx 60-70$% of the predicted values and do not depend on the size of $\delta\theta$; the values are more severely underestimated when smaller numbers of visibility periods are used, and the scatter about the average is also larger.
-Some of the astrometric excess noise being erroneously absorbed into other parts of the astrometric model is likely one contributing factor to the underestimation of $\delta\theta$.
-4.2 Empirical parallax and photocenter orbit semimajor axis error
-We construct an empirical error model for the parallax and photocenter orbit semimajor axis using the NSS catalog.
-The parallax error $\sigma_{\varpi}$ and photocenter orbit semimajor axis $\sigma_{a_{0}}$ are expected to be related to the centroiding error in the along-scan direction, $\sigma_{AL}$, and the number of visibility periods used, $N_{vis}$, according to
-$\sigma_{\varpi},\sigma_{a_{0}}\propto\frac{\sigma_{AL}(G)}{\sqrt{N_{vis}}}.$ (7)
-Using the number of visibility periods provides a more reliable estimate for position and proper motion uncertainties than the total number of good along-scan observations (Lindegren et al., 2018).
-To obtain an estimate for $N_{vis}$ in the population synthesis, we use the Gaia Observation Forecast Tool (GOST) [^7] https://gaia.esac.esa.int/gost/, which predicts when any given part of the sky crosses Gaia’s focal plane.
-To calculate $N_{vis}$, we evaluate the number of predicted scans at 49152 evenly spaced points on the sky, discard a random 10% of the predictions to mimic losses due to dead time [^8] The GOST page suggests 20% would be more appropriate, but we find 10% gives good agreement with observations., then calculate how many of the remaining predicted observations are separated by at least 4 days.
-We first attempted to find the constant of proportionality in Equation 7 relating $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ to $\sigma_{AL}(G)/\sqrt{N_{vis}}$.
-However, we find there is likely an additional unaccounted-for variable in the magnitude of the source (Appendix A).
-Thus, we instead rely on the uncertainties from the NSS catalog to make the estimates of $\sigma_{\varpi}$ and $\sigma_{a_{0}}$.
-In bins of 0.2 mag and integer values of visibility periods, we calculate the median and standard deviation $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in the NSS catalog.
-As expected, having more visibility periods and/or having a source magnitude $8\lesssim G\lesssim 13$ (where $\sigma_{AL}$ is smallest) tends to result in smaller values of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ (Figure 4).
-We find the standard deviation is roughly constant when the bin is well-sampled; we thus adopt a constant standard deviation of 0.12 mas for $\log_{10}\sigma_{\varpi}$ and 0.37 mas for $\log_{10}\sigma_{a_{0}}$.
-We also calculate the minimum value of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in each bin.
-Figure 4: Median parallax error (left) and photocenter orbit semimajor axis error (right) interpolated using the NSS catalog.
-The black dots show the location of the grid points in NSS; the inferred interpolated values are shown as the colored boxes.
-Note the color bar scale is different in the two panels.
-The median, standard deviation, and minimum values are used as the mean $\mu$, standard deviation $\sigma$, and minimum $a$, respectively, in a truncated normal distribution
-$f(x;\mu,\sigma,a,b)=\begin{dcases}\frac{1}{\sigma}\frac{\phi(x;\mu,\sigma)}{\Phi(a;\mu,\sigma)-\Phi(b;\mu,\sigma)},&a\leq x\leq b\\ 0,&\text{otherwise},\end{dcases}$ (8)
-where $\phi$ and $\Phi$ are the distributions of a normal PDF and a normal CDF, respectively.
-The truncation on the low end prevents artificially low errors; we do not apply the truncation on the higher end (hence the parameter $b$ in Equation 8 is infinite; in practice, it is set as a large number like 999).
-We then interpolate across this grid of values to determine the median $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ for a given magnitude and number of visibility periods used.
-Note that because this interpolation is performed over the NSS catalog, the interpolation is not defined for all pairs of ($N_{vis}$, $G$); the relationship would need to be extrapolated in order to estimate $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ at fainter magnitudes.
-The main concern with this approach is that by using the NSS catalog itself to estimate the errors, this will propagate whatever selection effects we are trying to characterize into the prediction itself.
-However, since the final forward model provides reasonable agreement with observations, we conclude that this approach is acceptable, if not ideal.
-4.3 Lack of sensitivity at $P\approx 1$ year and lack of solutions with $P>550$ days
-Figure 5: Empirical function that removes orbital periods around 1 year and at long orbital periods $P_{orb}\gtrsim 550$ days.
-This empirical filter is applied after the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters.
-The data used to make this figure are available through “Data behind the Figure”.
-The black outline in the topmost 1-D histogram in Figure 7 shows the distribution of all NSS astrometric orbital periods in Gaia.
-There is a clear dearth of solutions around 1 year because the astrometric signature of annual parallax is degenerate with $P_{orb}=1$ year.
-The 1-D histogram also shows a steep drop-off in the number of solutions at longer orbital periods.
-This is expected for $P_{orb}\gtrsim 1000$ days due to Gaia DR3 only containing 34 months of data.
-However, the number of solutions rapidly drops around $P_{orb}\approx 550$ days, which is much shorter than the Gaia DR3 temporal baseline.
-Forward modeling indicates that the ruwe  and parallax over parallax error filters do not cause the peak of the period distribution to be less than 2000 days (§5.1.2).
-We hypothesize that the drop-off in solutions at 550 days is primarily due to the astrometric processing cascade.
-As described in §2.3, acceleration solutions are tried before orbital solutions in the model cascade, and if an acceptable acceleration solution is found, it is kept and an orbital solution is not searched for, even if an orbital solution would have resulted in a better fit.
-Thus, long-period binaries are more likely to be satisfactorily fit with an acceleration model than binaries with orbital periods short enough to be covered multiple times over DR3.
-We test this hypothesis by cross-matching the nss_acceleration_astro  catalog to external binary catalogs to empirically estimate how many orbital solutions are lost to acceleration solutions in the model cascade (Appendix B).
-After accounting for the number of solutions at a given orbital period, we confirm that the number of acceleration solutions increases at longer periods, and the number of orbital solutions decreases at longer periods.
-This is also confirmed by the modeling by El-Badry et al.
-(2024) (their Figure 11).
-Modeling the dearth of solutions at $P_{orb}\approx 1$ year and the decrease in number of long $P_{orb}$ solutions require detailed knowledge of the spacecraft scanning law and individual epoch data.
-In the spirit of keeping calculations fast and analytic in this work, we construct a probabilistic function to empirically account for these effects.
-For $P_{orb}<250$ days, all solutions are kept; we assume that the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters fully capture the astrometric cascade.
-[^9] This is not entirely accurate; examination of the period distribution show that solutions are clearly lost around 190 days, which is 3×3$\times$3 × Gaia’s precession period.
-However, the number of solutions is so small we ignore this.
-We then assume the remaining discrepancy between the Gaia NSS astrometric orbits catalog and the forward model period distributions are due to the scanning law, finite Gaia DR3 temporal baseline, and loss of orbital solutions to acceleration solutions.
-We empirically determine this function by dividing the number of NSS orbital solutions by the number of solutions in the forward model after application of the ruwe  and $\varpi/\sigma_{\varpi}$ filters.
-We normalize the distribution assuming no solutions are lost at $P_{orb}=250$ days.
-Figure 5 shows the probabilistic function as a function of orbital period.
-We do not attempt to quantitatively understand the particular form of this function as this requires epoch-level modeling, which is outside the scope of this work.
-We also note that this empirical function is a function of $P_{orb}$ alone, which again is not completely accurate; for example, the 1 year orbit degeneracy can be more easily broken by highly eccentric orbits than circular orbits.
+For each $N_{vis}$ bin, a Gaussian is fit to the ratio $\delta\theta_{fit}/\delta\theta_{analytic}$.
+The mean of this Gaussian is taken to be the correction factor for that range of $P_{orb}$ and $N_{vis}$ (Figure 5).
+To determine the correction for a particular orbit, we interpolate across the $P_{orb}-N_{vis}$ grid to find the correction factor to $\delta\theta$, allowing the accurate prediction of ruwe  without actually having to generate and fit mock observations.
+We treat viewer-dependent quantities (e.g., argument of periastron, inclination, time of passage through periastron) as nuisance parameters; our probabilistic function is effectively integrating over these quantities.
+Although there are selection effects due to viewing angle (e.g., face-on orbits are more likely to show a larger astrometric signal than an edge-on orbit), these are not intrinsic binary properties.
+We are not concerned with characterizing these detection biases since these are randomized across all binaries.
+This enables us to understand the trends analytically, and the qualitative and quantitative effects of the fitting and modeling procedures.
+As shown in Figure 5, this correction is a strong function of orbital period.
+For $P_{orb}\gtrsim 1000$ days, the key assumption used to derive Equation 4 (having a fully-observed orbit) is violated; even for $P_{orb}\approx 1000$ days, this assumption may not hold if the orbit is not well-sampled for any reason (e.g., a highly eccentric orbit, or low number of visibility periods).
+At very long orbital periods $P_{orb}\gg 1000$ days, such a small fraction of the orbit will be covered by the Gaia DR3 temporal baseline that Equation 4 is no longer a valid approximation.
+The number of visibility periods used in the solution ($N_{vis}$) also affects the size of the erroneous attribution.
+For larger $N_{vis}$, the average size of the erroneous attribution is slightly smaller because more independent measurements better constrain the true orbit.
+We emphasize that we only model orbital periods shorter than 1500 days.
+This is approximately the range over which there is a still a useful correlation between the prediction of Equation 4 and the true ruwe.
+This choice is justified because in the NSS catalog, only 1% of solutions have $P_{orb}>1500$ days; this is a negligible fraction of solutions compared to the other uncertainties in the forward modeling.
+We recommend using the methodology of El-Badry et al.
+(2024) and working through the full framework of epoch modeling when trying to understand $P_{orb}>1500$ day solutions.
+Figure 5: Comparison of the astrometric excess noise $\delta\theta$ for sources in the synthetic catalog of El-Badry et al.
+(2024) as inferred by their ruwe, compared to the $\delta\theta$ as predicted by Equation 5.
+At longer orbital periods, the rescale factor gets smaller; when a orbit is only partially observed, it can be more easily be fit a single-star model, decreasing the astrometric excess noise.
+In addition, there are clear dips at orbital periods that are harmonics of 1 year or related to the scanning law.
+Finally, when the number of visibility periods is higher, the rescale factor is slightly larger (closer to unity) and the dips at the harmonics of 1 year are smaller.
+This is because when the orbit is better sampled, the solution is less prone to degeneracies that lead to incorrect solutions.
+4.3 Removal of acceleration solutions candidates
+Although understanding the selection effects of the acceleration solutions catalog is not the focus of this work, it must be modeled because it precedes the orbital solution modeling in the cascade.
+This is because a source that has an acceleration solution accepted in the main processing is removed from being a candidate for an orbital solution (Figure 1).
+As described in §2.3, acceleration solutions are tried before orbital solutions in the model cascade, and if an acceptable acceleration solution is found, an orbital solution is not searched for, even if an orbital solution would have resulted in a better fit.
+Thus, long-period binaries are more likely to be satisfactorily fit with an acceleration model than binaries with orbital periods short enough to be covered multiple times over DR3, and are likely to not be included in the orbit catalog.
+The intrinsic binary parameters that most strongly determine whether an acceleration solution is accepted are the orbital period $P_{orb}$ and the size of the photocenter orbit semimajor axis normalized by the measurement error $a_{0}/\sigma_{AL}$ [^13] Although a source’s positional uncertainty is determined by both its brightness and number of visibility periods, it is primarily a function of the brightness; thus, at the moment, we do not model the number of visibility periods..
+Eccentricity does not significantly affect the acceptance probability; see Appendix C for details.
+We use the simulated catalog of El-Badry et al.
+(2024) to construct an empirical probabilistic function to determine which sources will be candidates for orbital solutions.
+We start with all the NSS sources that pass the initial filtering step $(\texttt{ruwe}>1.4)$ and bin them by $P_{orb}$ and $a_{0}/\sigma_{AL}$.
+Within each $P_{orb}-a_{0}/\sigma_{AL}$ bin, we calculate the fraction of sources which are candidates for orbital solutions (sources where neither a 9-parameter or 7-parameter acceleration solution is accepted in the main processing).
+This is shown in Figure 6.
+To find the probability that a source with $\texttt{ruwe}>1.4$ is a candidate for an orbital solution, we interpolate across this $P_{orb}-a_{0}/\sigma_{AL}$ grid.
+The shapes of the high and low probability regions in Figure 6 can be understood as follows.
+A low probability of trying an orbital model is equivalent to a high probability of accepting an acceleration solution, and vice versa.
+Acceleration solutions are unlikely to be accepted at short orbital periods $P_{orb}\lesssim 500$ days because Gaia DR3 will have covered at least two full orbits, and an acceleration model will generally be a very poor fit; thus, the probability of trying an orbital solution is very high for $P_{orb}\lesssim 500$ days.
+At longer orbital periods $P_{orb}\gtrsim 500$ days, the probability of accepting an acceleration solution is maximal around the “sweet spot” of $\log_{10}(a_{0}/\sigma_{AL})\approx 0.8$ and $P_{orb}\approx 1300$ days.
+If $\log_{10}(a_{0}/\sigma_{AL})\lesssim 0.8$, then the detected acceleration will be noisy and less significant, making it less likely to pass the acceleration acceptance criterion on model significance.
+On the other hand, if $\log_{10}(a_{0}/\sigma_{AL})\gtrsim 0.8$, the astrometric signal will be much stronger, but the suitability of the acceleration model will depend on the orbital period.
+For fractionally covered orbits $P_{orb}\gtrsim 1000$ days, an acceleration model will be a good fit.
+However, for orbits $P_{orb}\approx 500-1000$ days where a full orbit or more has been covered, measurement noise is needed to mask the model mismatch.
+This explains the diagonal boundary separating the high and low probability regions for $P_{orb}\gtrsim 500$ days and $\log_{10}(a_{0}/\sigma_{AL})\gtrsim 0.8$ in Figure 6.
+Figure 6: Probability than an orbit model is tried for an NSS candidate, as a function of orbital period $P_{orb}$ and the photocenter orbit size $a_{0}$, normalized by the scan error $\sigma_{AL}$.
+Nearly all sources with orbital periods $P_{orb}\lesssim 500$ days are candidates for orbital solutions.
+However, for $P_{orb}\gtrsim 500$ days, the longer orbit solutions are increasingly unlikely to be candidates for an orbital solution, especially those with less significant photocenter orbit sizes.
+This is because these solutions are lost to the acceleration portion of the model cascade.
+As an independent test of whether or not longer period orbital solutions are preferentially assigned acceleration solutions, we cross-match the nss_acceleration_astro  catalog to external binary catalogs to empirically estimate how many orbital solutions are lost to acceleration solutions in the model cascade (Appendix D).
+After accounting for the number of solutions at a given orbital period, we confirm that the fraction of acceleration solutions increases at longer periods, and the fraction of orbital solutions decreases at longer periods.
+4.4 Orbital model acceptance criteria: parallax and photocenter semimajor axis error
+Figure 7: Parallax errors in the gaia_source (left) and NSS (right) catalogs.
+The analytic prediction (Equation 7) for the parallax errors and the true catalog errors are tightly correlated in the case of gaia_source (top panel); there is a tail of underestimated errors that we attribute to non-single sources being erroneously fit with single-star models (bottom panel).
+On the other hand, the analytic prediction and the true catalog errors are highly discrepant in the NSS catalog; the correlation is weak (top panel) and the analytic estimates systematically underestimate the errors (bottom panel).
+This is due to erroneous attribution of errors due to correlations between the parallax and orbit model parameters such as orbital period.
+This poor agreement necessitates an empirical correction to $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ in order to provide accurate uncertainties for NSS sources (Figure 8).
+In order to model the selection criteria of the orbital solutions, we need an estimate of the orbit significance $s_{orb}=a_{0}/\sigma_{a_{0}}$ and the parallax over error $\varpi/\sigma_{\varpi}$.
+The parameters $a_{0}$ and $\varpi$ are directly calculated from the binary source’s properties; we thus need a way to estimate their uncertainties.
+4.4.1 Analytic $\sigma_{\varpi},\sigma_{a_{0}}$ prediction
+The parallax error $\sigma_{\varpi}$ and photocenter orbit semimajor axis error $\sigma_{a_{0}}$ are expected to be proportional to the individual measurement errors, and decrease with the square root of the number of observations: $\sigma_{AL}(G)/\sqrt{N_{vis}}$.
+Using a random subset of the single-star fits in the gaia_source catalog, we find that the constant of proportionality is 0.48 for the parallax errors (Figure 7).
+Because we cannot measure the constant of proportionality for $\sigma_{a_{0}}$ using the gaia_source catalog since it consists of single-star fits, we indirectly infer it by inspecting the distribution of $\log_{10}(\sigma_{\varpi}/\sigma_{a_{0}})$ in the NSS catalog.
+It is symmetric with its peak at $\sigma_{\varpi}/\sigma_{a_{0}}\approx 1$, so we assume that the constant of proportionality for $\sigma_{a_{0}}$ is the same as for $\sigma_{\varpi}$.
+We thus adopt for our analytic error model
+$\sigma_{\varpi},\sigma_{a_{0}}=\frac{0.48\sigma_{AL}(G)}{\sqrt{N_{vis}}}.$ (7)
+Having the constant of proportionality exactly correct is not crucial, since we will be empirically rescaling it, as discussed in the next section.
+4.4.2 Empirical correction to analytic $\sigma_{\varpi},\sigma_{a_{0}}$ prediction
+In practice, Equation 7 will underestimate the true $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ for sources in the NSS astrometric orbits catalog (Figure 7).
+Analogous to the discussion in §4.2.2, this underestimation is due to the erroneous correlations between the parallax, position, proper motion and orbital parameters when performing the model fit.
+This erroneous error cannot be derived analytically, as it depends on the modeling of individual epoch observations.
+Thus, we construct an empirical $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ correction using the simulations of El-Badry et al.
+(2024), analogous to the empirical ruwe correction in §4.2.2.
+We analyze $3.3\times 10^{5}$ sources from the simulated cascade of El-Badry et al.
+(2024) that are fit with an orbital solution and that have reliably measured periods and eccentricities (i.e., the inferred $P_{orb}$ and $e$ are consistent within $3\sigma$ of the true $P_{orb}$ and $e$).
+Although the simulated cascade of El-Badry et al.
+(2024) fits $2.2\times 10^{6}$ orbital solution candidates, we only model the 15% of these sources with reliable fits as the remainder are spurious solutions that are not going to have accurate error estimates.
+Because El-Badry et al.
+(2024) are modeling the full Gaia DR3 cascade, they generate spurious solutions which are removed by the $F_{2}$ cut.
+These spurious solutions are not present in our analytic model, as they arise purely from the epoch modeling.
+We bin the reliably fit sources by their orbital periods and eccentricities.
+Within each bin, we fit a Gaussian to the distribution of $\log_{10}(\sigma_{\varpi,fit}/\sigma_{\varpi,analytic})$, where $\sigma_{\varpi,fit}$ is the parallax error as measured when modeling epoch astrometry in the catalog of El-Badry et al.
+(2024), and $\sigma_{\varpi,analytic}$ is the parallax error predicted from Equation 7.
+We then interpolate over this grid of values to sample the means and standard deviations, which are then sampled to assign a correction to $\sigma_{\varpi,analytic}$.
+We perform the same procedure on $\sigma_{a_{0},fit}$ and $\sigma_{a_{0},analytic}$ to determine a correction for the photocenter orbit semimajor axis $\sigma_{a_{0},analytic}$.
+The empirical correction is shown in Figure 8.
+The correction for $\sigma_{\varpi}$ is primarily a function of the orbital period.
+Equation 7 underpredicts $\sigma_{\varpi}$ by a factor of $10^{0.7}\approx 5$ for $P_{orb}\approx 1$ year, due to the degeneracy between $P_{orb}=1$ year orbits and parallax.
+The correction for $\sigma_{a_{0}}$ is both a function of eccentricity and orbital period.
+Solutions with either long orbital periods $P_{orb}\gtrsim 1000$ day or high eccentricities $e\gtrsim 0.6$ have $\sigma_{a_{0}}$ corrections around two orders of magnitude larger than solutions with lower eccentricity and shorter orbital period, excluding orbits with $P_{orb}\approx 1$ year.
+This demonstrates that the significance of the orbital solution is a function of the properties of the orbit itself.
+We emphasize that these empirical corrections to $\sigma_{a_{0}}$ and $\sigma_{\varpi}$ are only for solutions that have not received an acceleration solution in the NSS cascade, as they are derived only using sources that are fit with an orbital solution.
+We also emphasize that the corrections are specific to the observing strategy of Gaia DR3, and that they cannot be generally applied to observations that have a different temporal sampling, including Gaia DR4.
+Figure 8: Difference in predicted parallax error $\sigma_{\varpi}$ (left) and photocenter orbit semimajor axis $\sigma_{a_{0}}$ (right), as a function of orbital period, eccentricity, and photocenter orbit semimajor axis for a subset of bright, well-measured sources.
+The required rescaling of $\sigma_{\varpi}$ is primarily driven by orbital period, with it being significantly under-predicted at $P_{orb}\approx 1$ year, as well as being somewhat under-predicted at longer orbital periods ($P_{orb}\gtrsim 550$ days).
+The rescaling of $\sigma_{a_{0}}$ is more complex, with dependence on both orbital period and eccentricity.
+In addition to being under-predicted at $P_{orb}\approx 1$ year and $P_{orb}\gtrsim 1000$ day, it is also significantly underpredicted at high eccentricities ($e\gtrsim 0.6$).
 5 Results
@@
 The forward model begins with 150 million binaries within 2 kpc of the Sun at the present day.
-Even before applying the filters discussed in the previous section, many binaries are simply not detectable because they are too faint ($G\lesssim 21$) to be detected by Gaia, or their orbital periods are very long ($P_{orb}\gtrsim$ 30 years).
-In order to generate a more realistic estimate on the number of potentially detectable Gaia binaries, we calculate the number of binaries with $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which represents the range of orbital periods and magnitudes spanned in the NSS astrometric orbits catalog; 13 million binaries satisfy these criteria.
-Ultimately, after applying all the filtering cuts, we find $3.5\times 10^{5}$ binaries will be included in the NSS catalog.
-This is double the number of sources actually included in the NSS astrometric orbits catalog (168,065).
+Even before applying the filters discussed in the previous section, many binaries are simply not detectable because they are too faint ($G\gtrsim 21$) to be detected by Gaia, or their orbital periods are very long ($P_{orb}\gtrsim$ 30 years).
+In order to generate a more realistic estimate on the number of potentially detectable Gaia binaries, we calculate the number of binaries with $P_{orb}<1500$ days and $G<19$ mag, which represents the range of orbital periods and magnitudes spanned in the NSS astrometric orbits catalog; 12 million binaries satisfy these criteria.
+Ultimately, after applying all the filtering cuts, we find $2.6\times 10^{5}$ binaries will be included in the NSS catalog.
+This is about 50% more than the number of sources actually in the NSS astrometric orbits catalog (168,065).
 There are several possible explanations for this overestimate.
-One possibility is that the initial binary population produced by the binary population synthesis was incorrect in some way, e.g., the absolute number of systems was too high, the orbital period distribution of the binaries were incorrect and produced too many systems with $P_{orb}\approx 100-1000$ days, or the extinction was underestimated.
-Another is that the filtering steps do not accurately represent the model cascade, e.g., estimating the uncertainties $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ using the NSS astrometric catalog itself biases their values low.
+For example, the initial binary population produced by the binary population synthesis was incorrect in some way, e.g., the rescaling factor in §3 based on the GCNS was overestimated due to an overdensity of stars in the solar neighborhood, the orbital period distribution of the binaries was incorrect and produced too many systems with $P_{orb}\approx 100-1000$ days, the extinction was underestimated, or some binaries were disrupted by stellar encounters or cluster dispersal.
 Nonetheless, the forward model is able to reproduce the total number of sources within a factor of two, which will still provide useful constraints in population studies.
-Table 1: NSS filters and resultant number of predicted binaries
-Filter No. of binaries (%)
-None (all binaries) $1.5~{}\times~{}10^{8}$ (100%)
-$15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag $1.3~{}\times~{}10^{7}$ (8.8%)
-ruwe$>1.4$ $6.3~{}\times~{}10^{6}$ (4.2%)
-$\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ $1.2~{}\times~{}10^{7}$ (8.2%)
-$a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{ day}})$ $8.6~{}\times~{}10^{6}$ (5.8%)
-ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ $5.6~{}\times~{}10^{6}$ (3.7%)
-ruwe$>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$, $P_{orb}\approx$ 1 yr, $P_{orb}>550$ day $3.5~{}\times~{}10^{5}$ (0.2%)
+Table 1: Selection effect of NSS cascade filters
+Filter No. of binaries (%) Selection effect
+None (all binaries) $1.5~{}\times~{}10^{8}$ (100%) N/A
+NSS potentially detectable ($P_{orb}<1500$ day, $G<19$ mag) $1.2~{}\times~{}10^{7}$ (8%) Preferential selection of primaries with $M_{1}\approx 1M_{\odot}$
+NSS catalog candidates: ruwe$>1.4$ $1.3\times 10^{6}$ (0.9%) Preferentially selects $P_{orb}\approx 10^{3\pm 1}$ days
+Orbital solution candidates (i.e., not selected to acceleration solution candidates): $s_{7,9}<12$ or $\omega/\sigma_{\omega}>2.1s_{7,9}^{1.05}$ $7.4\times 10^{5}$ (0.5%) Preferentially removes $P_{orb}\gtrsim 500$ days, increasing from 0% at $P_{orb}\approx 500$ days up to 85% at $P_{orb}\approx 2000$ days
+Orbital solutions ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}}$) $2.6\times 10^{5}$ (0.2%) Preferentially removes 1) high eccentricity $e\gtrsim 0.6$ solutions; 2) short-period $P_{orb}\lesssim 500$ days; 3) $P_{orb}\approx 1$ year solutions
 Note.
-— The first line, where Filter is “None”, denotes the number of binaries within 2 kpc of the Sun.
-The filter on $15<P_{orb}<10^{4}$ day, $18.54<G<$ 18.5 mag is not an actual filter criterion for the NSS astrometric orbits catalog, but is a more representative number of the possible number of orbital systems that could be detected by Gaia as compared to all binaries within 2 kpc, which has many systems with orbital periods too long or magnitudes too faint to be detected.
+— Steps in the NSS astrometric processing cascade, the resulting number of binaries, and a summary of the selection effect introduced by the filter.
+Note that “NSS potentially detectable” is not a filter described in Halbwachs et al.
+(2023).
+It is included as a proxy for the fraction of the total binary population that could realistically receive an astrometric solution; most binaries are too faint to be detected by Gaia or have orbital periods significantly longer than the $\approx 3$ year time baseline of DR3.
 5.1.2 Distribution of binary and orbital parameters
-Figure 6: Effect of filtering on $\texttt{ruwe}>1.4$ (ruwe), $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ ($\varpi/\sigma_{\varpi}$), $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ ($a_{0}/\sigma_{a_{0}})$, and $P_{orb}\approx 1$ yr and $P_{orb}>550$ day ($P$) on the period distribution in the forward modeling.
-Figure 7: 1-D and 2-D PDFs in the period $P_{orb}$, parallax $\varpi$, eccentricity $e$, photocenter orbit semimajor axis $a_{0}$, ruwe, apparent magnitude $G$, and parallax over parallax error $\varpi/\sigma_{\varpi}$ of the forward model (red) as compared to the actual NSS astrometric orbits catalog (black).
-In general, the forward modeling does a good to reasonable job in matching the actual NSS distributions including the correlations, with the exception of the eccentricity distribution.
 Figure 9: Effect of filtering on the distribution of orbital period $P_{orb}$, parallax $\varpi$, eccentricity $e$, ruwe, magnitude $G$, primary and secondary mass $M_{1},M_{2}$, luminosity ratio $l_{G}$, mass ratio $q$, and photocenter scale $\delta_{ql}$.
-The dotted line shows the birth distribution from Moe & Di Stefano (2017).
-The black outline shows the distribution of systems that have orbital periods between $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia.
-The filter on long-period orbits preferentially removes larger photocenters and large ruwe.
+The dashed gray line shows the birth distribution from Moe & Di Stefano (2017).
+The dotted gray line shows the distribution of systems that have orbital periods $P_{orb}<1500$ days and $G<19$ mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia.
 The magnitude cut preferentially removes distant and low-mass binaries.
 It causes the primary mass distribution to peak roughly around $0.8M_{\odot}$ (when binned logarithmically).
-The gray outline shows the systems that satisfy $\texttt{ruwe}>1.4$.
-Finally, the red outline shows the systems that satisfy all the filtering criteria applied ($\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$, $P_{orb}\approx 1$ year, $P_{orb}>550$ day).
-Figure 6 summarizes the effect of the filtering steps on the period distribution.
-The $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ filters all preferentially remove shorter period systems.
-The ruwe  filter removes more systems overall, while $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$ filters remove shorter period solutions more strongly.
+The solid black line shows sources that are NSS model candidates ($\texttt{ruwe}>1.4$), the blue line shows sources that are orbital model candidates (i.e., NSS model candidates that are not candidates for an acceleration solution), and the red line shows the sources that are in the final astrometric orbits catalog.
+Zoom-ins of the $P_{orb}$ and $e$ panels are shown in Figure 10.
+Figure 10: Effect of the different filters in the NSS processing cascade on the distribution of orbital periods (left) and eccentricity (right).
+Each line shows a different filter: $\texttt{ruwe}>1.4$ (black dotted), orbit candidates (gray dashed), and the final orbit catalog (red solid).
+The filters in the acceleration cascade preferentially remove long period orbital solutions.
+In the orbital cascade, the filter on $\varpi/\sigma_{\varpi}>20,000\textrm{ d}/P_{orb}$ preferentially removes short period and 1-year period solutions, while the filter on $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$ also preferentially removes eccentric solutions.
 Figure 2 is a corner plot comparing the orbital period, parallax, eccentricity, photocenter orbit semimajor axis, ruwe, apparent magnitude, and parallax over parallax error of the mock forward model to the actual NSS astrometric orbit catalog.
-The 2-D correlations are broadly preserved, and the 1-D distributions are also generally reasonable matches.
-The deficit of solutions around $P_{orb}\approx 1$ year and the decline in number of solutions longer than $P_{orb}>550$ days in the forward modeled period distribution were tuned to match the NSS catalog, so they are unsurprisingly a reasonable match.
-The parallax, magnitude, and parallax over parallax error distributions are also a very good match.
-The photocenter orbit semimajor axis and ruwe  distributions are also reasonable matches, although the models tend to result in distributions slightly more sharply peaked than the NSS catalog.
-The eccentricity distribution between the model and the true catalog are by far the most discrepant, with the Gaia NSS distribution peaking around $e\approx 0.2$, while the forward modeled distribution peaks at a much higher eccentricity of $e\approx 0.75$.
-We discuss possible reasons for this in §6.1.
-Although the parallax error and photocenter orbit semimajor axis errors were interpolated directly using the NSS catalog itself, the filters and distributions involving $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ were not tuned post-facto like the period distribution.
-Figure 8 compares the 1-D distribution of binary parameters in the initial birth population with that obtained from a simple filter on observability, which we define to be $15<P_{orb}<10^{4}$ days and $18.54<G<$ 18.5 mag, which is the range spanned by the NSS data and a rough proxy for systems that could be characterized by Gaia.
-It also shows the distribution of parameters after application of the $\texttt{ruwe}>1.4$ filter and all NSS filters.
-As discussed previously, the ruwe  filter trims short orbital period, small $a_{0}$, distant, and faint systems.
-Including the rest of the filters does not significantly change the 1-D distribution of the parameters as compared to the $\texttt{ruwe}>1.4$ filter, except the orbital period distribution.
+The 2-D correlations are preserved, and the 1-D distributions are also generally reasonable matches.
+The orbital period distribution is a good match between the forward model and actual NSS catalog.
+The distribution of parallaxes in the forward model are systematically offset larger than the distribution of NSS parallaxes by about 0.1 dex; otherwise, the agreement is also good.
+The distribution of $a_{0}$ and ruwe are excellent matches.
+The forward model apparent magnitude distribution is somewhat less peaked and has slightly more bright sources than the NSS distribution, but the match is otherwise good.
+The distribution of parallax over error in the forward model are systematically offset larger than the distribution of NSS parallaxes by about 0.3 dex; part of this may be related to the parallaxes being high by 0.1 dex.
+The eccentricity distribution of the model and the true catalog are the most discrepant parameter, with the Gaia NSS distribution peaking around $e\approx 0.2$, while the forward modeled distribution is flatter and has a sharp drop-off in the number of solutions around $e\approx 0.6$; we discuss possible reasons for this in §6.1.
+5.2 Cascade selection effects
+Next, we explore the selection effects of each part of the cascade and how they produce the resultant distributions in Figure 2, in particular for the orbital period and eccentricity distributions.
+Figure 9 compares the 1-D distribution of binary parameters in the initial birth population (dashed gray line) with that obtained from a simple filter on systems that could be astrometrically characterized by Gaia (dotted gray line), defined to be $P_{orb}<1500$ days and $G<19$ mag.
+It also shows the distribution of parameters after application of the $\texttt{ruwe}>1.4$ filter (black line), removal of acceleration solutions (blue line), and the final NSS astrometric orbit catalog (red line).
+Table 1 summarizes the effects of each filter on the population of sources.
+We now discuss in detail the impact of each step in the NSS astrometric cascade.
+5.2.1 Input catalog
+The initial selection of NSS astrometric candidates primarily introduces a selection effect on the orbital period of the solutions.
+The $\texttt{ruwe}>1.4$ filter preferentially selects sources with $P_{orb}\approx 1000$ days and the orbital period distribution is roughly symmetric in $\log_{10}(P_{orb})$ (Figure 9).
+As previously mentioned, our ruwe model is only accurate for $P_{orb}<1500$ days, so we validated that the decrease in number of solutions held out to $P_{orb}=10,000$ days using the catalog of El-Badry et al.
+(2024).
+Orbits with $P_{orb}\approx 10^{3\pm 1}$ days are the sweet spot in the ruwe criteria: those with $P_{orb}<100$ days will be too small to have a significant photocenter wobble given the measurement uncertainties; those with $P_{orb}>10,000$ days will lack significant orbital coverage during Gaia DR3 and can be approximately modeled by a single-star fit, given the measurement uncertainties.
+An orbit of 1000 days maximizes the size of the photocenter wobble and orbital coverage during the survey.
+Next, we consider the effects on the eccentricity distribution (Figure 9).
+Short period orbits are preferentially removed by the ruwe  filter, and most short period orbits $P_{orb}\lesssim 30$ days are circular (Moe & Di Stefano, 2017).
+Although ruwe  also directly removes high eccentricity orbits (Equation 6), it is a weak effect.
+Finally, the ruwe filter also preferentially selects nearby sources with large photocenters and photocenter scales (Figure 9).
+5.2.2 Acceleration models
+Next, we analyze the effect of the acceleration models. Recall that sources that have an accepted acceleration model during the main NSS processing are not able to be orbital solution candidates.
+The acceleration portion of the NSS cascade preferentially removes solutions with $P_{orb}>500$ days (Figure 10).
+Between $P_{orb}\approx 500-2000$ days, the fraction of NSS astrometric candidates that have accepted acceleration solutions increases from 0 to about 85%.
+Because long $P_{orb}$ solutions are more likely to only have partial orbital period coverage, the acceleration models are more likely to be accepted; orbits with full coverage are more likely to be poor fits to the acceleration models.
+The acceleration portion of the NSS cascade explains why the peak of the orbital period distribution is at such a short $P_{orb}\approx 500$ days.
+In the absence of the acceleration cascade, one would expect high sensitivity to $P_{orb}\approx 700-900$ day sources given the $\approx 1000$ day time baseline of Gaia DR3.
+The removal of acceleration solutions does not substantially change the eccentricity distribution (Figure 10).
+5.2.3 Orbital model selection criteria
+Finally, the last step is selecting the orbital model candidates that will actually be accepted into the final catalog.
+The orbital model significance filter $a_{0}/\sigma_{a_{0}}>5$ has a very strong effect on the eccentricity distribution, preferentially removing eccentric solutions $e\gtrsim 0.6$.
+It also preferentially removes long-period solutions $P_{orb}\gtrsim 1000$ days.
+This is because at a fixed source magnitude and number of visibility periods, solutions which do not have good orbital coverage will have larger uncertainties on $a_{0}$.
+Similarly, eccentric solutions will also have larger uncertainties on $a_{0}$.
+The $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{d}}$ filter also preferentially removes eccentric systems, although it preferentially removes short period systems as well because of the $1/\sqrt{P_{orb}}$ dependence.
+The parallax significance relative to orbital period filter $\varpi/\sigma_{\varpi}>20,000\textrm{d}/P_{orb}$ does not bias against the eccentricity.
+It preferentially removes short period solutions due to the $1/P_{orb}$ dependence.
+It also preferentially removes solutions with $P_{orb}\approx 1$ year because these sources have a very low $\varpi/\sigma_{\varpi}$, due to the large errors from the degeneracy between parallax and orbital motion.
+Although this effect can be seen in the previous two filtering steps, it is much stronger in this step.
 5.3 Probability of finding Gaia black holes
@@
 Name RA (deg) Dec (deg) $P_{orb}$ (day) $e$ $\varpi$ (mas) $M_{*}(M_{\odot})$ $M_{\bullet}(M_{\odot})$ $G_{*}$ (mag) $M_{G,0*}$ (mag) $A_{G}$ (mag) p(NSS)
-Gaia BH1 262.17 –0.58 186 0.45 2.09 0.93 9.62 13.77 4.6 0.76 0.33
-Gaia BH2 207.57 –59.24 1277 0.52 0.859 1.07 8.94 12.28 1.8 0.15 0.04
-Gaia BH3 294.83 14.93 4195 0.73 1.67 0.76 32.7 11.23 2.12 0.22 0.00
+Gaia BH1 262.17 –0.58 185 0.43 2.09 0.93 9.24 13.77 4.6 0.76 0.38
+Gaia BH2 207.57 –59.24 1277 0.52 0.859 1.07 8.94 12.28 1.8 0.15 0.27
 Note.
-— Properties of Gaia BH1, BH2, and BH3.
+— Properties of Gaia BH1 and BH2.
 * subscript denotes properties of the stellar companion, $\bullet$ subscript denotes properties of the black hole.
-All values except $M_{G,*}$ and $A_{G}$ are taken from El-Badry et al.
-(2023a), El-Badry et al.
-(2023b), and Gaia Collaboration et al.
-(2024) for Gaia BH1, BH2, and BH3, respectively.
+All values except $M_{G,*}$ and $A_{G}$ are taken from Nagarajan et al.
+(2024) and El-Badry et al.
+(2023b) for Gaia BH1 and BH2, respectively.
 We note for Gaia BH2 that our extinction map suggests much less extinction ($\approx 0.4$ mag) than the map used by El-Badry et al.
 (2023b) and so our inferred extinction-corrected absolute magnitude of the star $M_{G,0*}$ in Gaia BH2 is fainter.
-Similarly for Gaia BH3 our extinction map suggests $\approx 0.3$ mag less extinction than Gaia Collaboration et al.
-(2024).
-We apply our forward modeling framework to determine the probability of inclusion of Gaia BH1 (El-Badry et al., 2023a; Chakrabarti et al., 2023), Gaia BH2 (El-Badry et al., 2023b), and Gaia BH3 (Gaia Collaboration et al., 2024) in the Gaia DR3 NSS astrometric orbits catalog.
+We do not calculate the probability for Gaia BH3 Gaia Collaboration et al.
+(2024) because its orbital period is longer than 1500 days.
+We apply our forward modeling framework to determine the probability of inclusion of Gaia BH1 (El-Badry et al., 2023a; Chakrabarti et al., 2023) and Gaia BH2 (El-Badry et al., 2023b) in the Gaia DR3 NSS astrometric orbits catalog.
+We do not calculate the probability for Gaia BH3 (Gaia Collaboration et al., 2024) because its orbital period is much longer than 1500 days, where our ruwe  model breaks down.
 We calculate the probability as a function of sky position, orbital period, eccentricity, parallax, and the component magnitudes and masses.
-We integrate over random inclinations in $|\cos i|$.
-We find the probability of Gaia BH1 ((RA, Dec) = (262.17, –0.58) deg, $P_{orb}=186$ day, $e=0.45$, $\varpi=2.09$ mas, $M_{*}=0.93M_{\odot}$, $M_{\bullet}=9.62M_{\odot}$, $G_{*}=13.77$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 33%.
+We integrate over random inclinations that are uniform in $|\cos i|$.
+We find the probability of Gaia BH1 ((RA, Dec) = (262.17, –0.58) deg, $P_{orb}=186$ day, $e=0.45$, $\varpi=2.09$ mas, $M_{*}=0.93M_{\odot}$, $M_{\bullet}=9.62M_{\odot}$, $G_{*}=13.77$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 38%.
 We also vary the location of BH1 on the sky and its distance from the Sun (Figure 11) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog changes.
-Since Gaia BH1 was located near the ecliptic plane, it had a relatively low probability of inclusion; toward the ecliptic poles, the inclusion probability would have been closer to 70–90%.
-Gaia BH1 was also relatively close by, at a distance of about 500 pc.
+Since Gaia BH1 is located near the ecliptic plane, it had a relatively low probability of inclusion; toward the ecliptic poles, the inclusion probability would have been closer to 60–80%.
+Gaia BH1 is also relatively close by, at a distance of about 500 pc.
 Had it been at the same (RA, Dec) but slightly farther away, at 600 pc or more, it would not have made it into the NSS catalog.
 Conversely, had it been slightly closer, 400 pc or less, it would have had $>80\%$ probability of making it into the catalog.
-This suggests that if there were Gaia BH1-like systems within a few hundred pc of the Sun, they should have been included in the Gaia DR3 NSS astrometric orbits catalog; the lack of such detections suggests that there are no or few Gaia BH1-like sources within 300–400 pc of the Sun.
-We find the probability of Gaia BH2 ((RA, Dec) = (207.57, –59.24), $P_{orb}=1277$ day, $e=0.52$, $\varpi=0.86$ mas, $M_{*}=1.07M_{\odot}$, $M_{\bullet}=8.94M_{\odot}$, $G_{*}=12.28$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 4%.
+This suggests that if there are Gaia BH1-like systems within a few hundred pc of the Sun, they should have been included in the Gaia DR3 NSS astrometric orbits catalog; the lack of such detections suggests that there are no or few Gaia BH1-like sources within 300–400 pc of the Sun.
+We find the probability of Gaia BH2 ((RA, Dec) = (207.57, –59.24), $P_{orb}=1277$ day, $e=0.52$, $\varpi=0.86$ mas, $M_{*}=1.07M_{\odot}$, $M_{\bullet}=8.94M_{\odot}$, $G_{*}=12.28$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 27%.
 We also vary the location of BH2 on the sky and its distance from the Sun (Figure 11) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog varies.
-Since Gaia BH2 has a long orbital period, the main factor of whether it made it into the NSS catalog is whether or not it made it through the astrometric cascade or not.
-The inclusion probability of Gaia BH2 does not change significantly with distance until beyond $\gtrsim 2.5$ kpc, at which point the inclusion probability substantially drops.
-The inclusion probability also does not vary significantly across the sky.
-This suggests that the main factor of whether Gaia BH2 made it through the cascade or not is the long period filtering.
-The probability of an orbital solution being kept at $P_{orb}=1277$ days is 4% (Figure 5).
-This suggests that Gaia BH2 was a relatively lucky detection and that there are many more such systems that will be detectable in Gaia DR4, either in an updated NSS catalog, or in the epoch astrometry.
-We find the probability of Gaia BH3 to be included in the Gaia DR3 NSS astrometric catalog to be 0%.
-This is because of the orbital period being $>4000$ days; none of these systems pass the long period filtering.
-This is not entirely correct, as $<0.5\%$ of solutions in the NSS astrometric orbits catalog have $P_{orb}>2000$ days, and $<0.05\%$ of solutions have $P_{orb}>4000$ days, so it is possible for such long solutions to be in the catalog.
-However, within a margin of a few percent, this is probably accurate.
-In addition, Gaia BH3 was not included in any of the NSS catalogs, orbital (nss_two_body_orbit), nor acceleration (nss_acceleration_astro, nss_non_linear_spectro).
+Since Gaia BH2 has a long orbital period, the main factor of whether it made it into the NSS catalog is whether or not it was assigned an acceleration solution or not.
+The inclusion probability shows a weak dependence on ecliptic latitude, and does not vary significantly ($<10$%) across the sky.
+Gaia BH2 is around 1.2 kpc from the Sun.
+Had it been at the same (RA, Dec) but beyond 3 kpc, it would not have been included in the NSS catalog.
+Conversely, had it been closer (between 500 and 1000 pc), it would have a somewhat higher probability ($50-60$%) of making it into the NSS catalog.
+The results here for BH1 are consistent with the results in El-Badry et al.
+(2024), which find a 39% probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog.
+The probability of inclusion for BH1 as a function of distance (Figure 15 of El-Badry et al.
+2024) is also very similar.
+For BH2, El-Badry et al.
+(2024) finds a higher probability of 48%.
+The probability of inclusion for BH2 as a function of distance (Figure 15 of El-Badry et al.
+2024) is also similar, albeit their inclusion probabilities are generally about $10-20$% higher.
+Since our model was tuned to match that of El-Badry et al.
+(2024), we might expect the results to be more similar.
+This may be due to the probabilistic nature of the acceleration cascade modeling; $P_{orb}$ and $a_{0}/\sigma_{AL}$ may not capture the full dependencies of probability acceptance, leading to large scatter.
+This suggests the methodology may be reliable only to a factor of ${\sim}2$ beyond $P_{orb}\gtrsim 500$ days.
 Figure 11: Top panels: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of sky position (192 equally spaced points on sky); the color of the circle denotes the probability of inclusion.
@@
 For Gaia BH1, the detectability as a function of sky position clearly shows the pattern of the scanning law.
-For Gaia BH2, the detectability does not have a dependence on the sky position.
+The ecliptic plane receives fewer scans than the ecliptic poles (Lindegren et al., 2012), reducing the astrometric precision and in turn the size of a detectable astrometric orbit.
+Gaia BH1 thus tends to be more reliably detected around the ecliptic poles than the ecliptic plane.
+For Gaia BH2, the detectability does not have as strong a dependence on the sky position.
 Bottom panel: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of the system’s distance.
@@
 For Gaia BH1-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+4.6$.
-We consider orbital periods $200010<P_{orb}<$ 2000 days, spanning roughly the range of detected NSS astrometric orbits.
+We consider orbital periods $150030<P_{orb}<$ 1500 days, spanning roughly the range of detected NSS astrometric orbits.
 We will refer to these as Sun-like star + BH binaries, or Sun+BH binaries for short.
 For Gaia BH2-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+1.8$.
-We consider orbital periods $2000300<P_{orb}<$ 2000 days.
+We consider orbital periods $1500300<P_{orb}<$ 1500 days.
 The minimum orbital period is 300 days such that the semimajor axis of the orbit is larger than the typical red giant radius.
@@
 detected systems.
-On average, 3.5% of Sun+BH binaries with orbital periods $200010<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable, and 16% of RG+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable [^10] On average, 7.0% of Sun+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable..
+On average, 5% of Sun+BH binaries with orbital periods $150030<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable, and 27% of RG+BH binaries with orbital periods $1500300<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable [^14] On average, 10% of Sun+BH binaries with orbital periods $1500300<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable..
 The period distribution unsurprisingly looks similar to the NSS astrometric orbits catalog.
@@
 RG+BH systems can be detected to at least 2 kpc because the red giant is much more luminous than a Sun-like star despite having the same mass.
-There is no strong dependence of the detectability of the system on black hole mass.
+There is no strong dependence of the detectability of the system on black hole mass within the $5-80M_{\odot}$ range.
 The size of the photocenter orbit semimajor axis $a_{0}$ (Equation 1) increases by a factor of only 2.4 for a $5M_{\odot}$ vs.
@@
 We can use this model to make a simple argument on the occurrence rate of detached black holes in the Milky Way in various orbital period ranges.
-We first consider Sun+BH binaries with orbital periods $400100<P_{orb}<$ 400 days, $1000400<P_{orb}<$ 1000 days, and $20001000<P_{orb}<$ 2000 days, and separately RG+BH binaries with orbital periods $1000400<P_{orb}<$ 1000 days and $20001000<P_{orb}<$ 2000 days.
-We assume that Gaia BH1 (Sun+BH with an orbital period $400100<P_{orb}<$ 400 days) and BH2 (RG+BH with an orbital period $20001000<P_{orb}<$ 2000 days) are the only two star + BH binaries in the Gaia DR3 astrometric orbits catalog.
-We then use our model to calculate the number of detectable detached BHs within 1 kpc of the Sun within these period ranges.
+We first consider Sun+BH binaries with orbital periods $400100<P_{orb}<$ 400 days, $1000400<P_{orb}<$ 1000 days, and $15001000<P_{orb}<$ 1500 days, and separately RG+BH binaries with orbital periods $1000400<P_{orb}<$ 1000 days and $15001000<P_{orb}<$ 1500 days.
+We assume that Gaia BH1 (Sun+BH with an orbital period $400100<P_{orb}<$ 400 days) and BH2 (RG+BH with an orbital period $15001000<P_{orb}<$ 1500 days) are the only two star + BH binaries in the Gaia DR3 astrometric orbits catalog, as the studies that have identified candidates in the astrometric NSS catalog (Andrews et al., 2022; Shahaf et al., 2023) have been followed up and not yielded detections.
+Additional candidates identified in the Gaia BH1 discovery papers (El-Badry et al., 2023a; Chakrabarti et al., 2023) also have been ruled out.
+We then use our model to calculate the number of detectable detached BHs within 2 kpc of the Sun within these period ranges.
 For the period ranges where no BH has been detected, we take this to be the upper limit on the number of those BH binaries.
-Assuming there are $\approx$100-200 billion stars in the Galaxy, 160 million, or 0.1% of those stars are within 1 kpc of the Sun; to convert our number of objects to within 1 kpc to the number of objects in the Galaxy, we multiply by 1000.
+Assuming there are 100 billion stars in the Galaxy, 1 billion, or 1% of those stars are within 2 kpc of the Sun; to convert our number of objects within 2 kpc to the number of objects in the Galaxy, we multiply by 100.
 Table 3 lists the detection probabilities.
-We find that the Galaxy hosts $5000\pm 2000$ Sun+BH binaries with $400100<P_{orb}<$ 400 days, and $20,000\pm 4000$ RG+BH binaries with $20001000<P_{orb}<$ 2000 days.
-We also find that there are at most 2000 Sun+BH binaries with $1000400<P_{orb}<$ 1000 day, 20,000 Sun+BH binaries with $20001000<P_{orb}<$ 2000 day, and 20,000 RG+BH binaries with $20001000<P_{orb}<$ 2000 day.
-Note the limits and constraints are weaker for the long-period end, due to the sharp decrease in number of solutions beyond 1000 days.
+We find that the Galaxy hosts $3000\pm 500$ Sun+BH binaries with $400100<P_{orb}<$ 400 days, and $400\pm 200$ RG+BH binaries with $15001000<P_{orb}<$ 1500 days.
+We also find that there are $<800$ Sun+BH binaries with $1000400<P_{orb}<$ 1000 day, $<12,000$ Sun+BH binaries with $15001000<P_{orb}<$ 1500 day, and $<300$ RG+BH binaries with $1000400<P_{orb}<$ 1000 day.
+These uncertainties are assuming Poisson statistics; they do not account for systematic errors (e.g., errors in the forward model, or the assumed number of stars in the Galaxy.)
 Although the previous paragraphs treats the constraints on Sun+BHs separately from RG+BHs separately, in reality these constraints are related, as Sun-like stars will eventually evolve into red giants.
 Because the evolutionary timescale of the red giant phase is only a $\sim 10\%$ of a star’s main sequence lifetime, the number of RG+BHs should also be a few percent of the number of Sun+BH binaries in a given period range.
-This is in tension with the results for the $20001000<P_{orb}<$ 2000 range, as our model predicts more RG+BH binaries than Sun+BH binaries.
-This is because the empirical filter that removes long-period solutions probabilistically places an upper limit on the number of solutions.
-For sources that pass the $\texttt{ruwe}\ >1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ days}/P_{orb}$ filters, this limit is $\sim 4\%$ for $20001000<P_{orb}<$ 2000.
-Because of this, the number of RG+BHs in Table 3 is probably an overestimate.
+This is generally consistent with the results we find.
+For the $1000400<P_{orb}<$ 1000 day range, we only have upper limits for RG+BH and Sun+BH binaries and so this is not particularly constraining.
+For the $15001000<P_{orb}<$ 1500 day range, this implies that the number of RG+BH binaries are likely on the lower side of the estimated range, and for the Sun+BH binaries, the number of systems will be close to the upper limit and they just barely managed to escape inclusion in the NSS astrometric orbits catalog.
 In comparison, El-Badry et al.
@@
 Table 3: Number of black holes
-$P_{orb}$ range (day) Prob. detect N($d<1$ kpc)
+$P_{orb}$ range (day) Prob. detect N($d<2$ kpc)
 Sun-like star + BH
-$400100<P_{orb}<$ 400 0.18 $5\pm 2$
-$1000400<P_{orb}<$ 1000 0.44 $<2$
-$20001000<P_{orb}<$ 2000 0.05 $<20$
+$400100<P_{orb}<$ 400 0.03 $30\pm 5$
+$1000400<P_{orb}<$ 1000 0.14 $<8$
+$15001000<P_{orb}<$ 1500 0.09 $<12$
 Red giant + BH
-$1000400<P_{orb}<$ 1000 0.50 $<2$
-$20001000<P_{orb}<$ 2000 0.05 $20\pm 4$
+$1000400<P_{orb}<$ 1000 0.36 $<3$
+$15001000<P_{orb}<$ 1500 0.25 $4\pm 2$
 Note.
-— Number of black holes within 1 kpc of the Sun.
-Upper limits and uncertainties are $1\sigma$.
-To convert these results to the number of black holes in the entire Milky Way, multiply the values in the N($d<1$ kpc) column by 1000.
+— Number of black holes within 2 kpc of the Sun.
+Upper limits and uncertainties are $1\sigma$, assuming Poisson errors.
+To convert these results to the number of black holes in the entire Milky Way, multiply the values in the N($d<2$ kpc) column by 100.
 6 Discussion
 6.1 Discrepancy in eccentricity
-Figure 11: Top panel: eccentricity distribution correction factor derived from the forward modeling.
-Bottom panel: NSS eccentricity distribution for Sun-like stars with orbital periods $1000100<P_{orb}<$ 1000 days, before and after correction.
-Note that we only plot orbits with $e>0.01$.
-The blue line shows a fit to the corrected distribution: a beta function with $\alpha=1.7$, $\beta=3.2$, rescaled with a minimum of 0.017 and width of 0.98.
-The ruwe  filter does not have a strong bias against eccentricity, so the raw and corrected distributions look very similar.
-There is a serious discrepancy between the Gaia NSS astrometric orbit and forward modeled eccentricity distribution, with the forward model predicting significantly more eccentric binaries than are present in the Gaia astrometric orbits catalog.
+There is a discrepancy between the Gaia NSS astrometric orbit and forward modeled eccentricity distribution, with the forward model predicting significantly more eccentric binaries than are present in the Gaia astrometric orbits catalog.
 We hypothesize that the starting eccentricity distribution used in the forward model is incorrect and leads to this discrepancy.
@@
 This parametrization results in $\eta=0.13$ for $P_{orb}=100$ days and $\eta=0.32$ for $P_{orb}=1000$ days, and was the assumed eccentricity distribution used in the forward model presented in this work.
-We use our forward model to determine a correction that can be applied to the observed NSS eccentricity distribution to convert it into the true underlying eccentricity distribution.
-We calculate the ratio of the normalized distribution of mock NSS astrometric orbits solution to the normalized distribution of the birth population as a function of eccentricity in four period bins.
-The reason for binning the results by period is that the eccentricity distribution is a function of period in the forward model (Equation 9).
-We then average the ratios across the period bins, and fit a quadratic through this average to determine a correction factor on the eccentricity distribution as a function of eccentricity (Figure 11).
-As expected, the astrometric cascade has a slight preference for filtering out high-eccentricity systems.
-This is due to the $\texttt{ruwe}>1.4$ filter (§4.1.1).
-El-Badry et al.
-(2024) find a similar discrepancy between the Gaia NSS astrometric orbit and their forward modeled eccentricity distribution, with their forward model also predicting more eccentric binaries than are present in the Gaia astrometric orbits catalog.
-However, their discrepancy is not as significant as found in this work; they find their astrometric cascade produces a significantly stronger bias against the detection of eccentric orbits than we do.
-This bias is due to the eccentric orbits having larger astrometric uncertainties $\sigma_{a_{0}}$ because they are preferentially sampled near apastron; these solutions are then removed by the $a_{0}/\sigma_{a_{0}}$ filter.
-Although we do model the $a_{0}/\sigma_{a_{0}}$ filter in this work, our model for $\sigma_{a_{0}}$ is only a function of the number of visibility periods $n_{vis}$ and magnitude $G$ and does not account for the eccentricity or period of the orbit (§4.2).
-Thus, it does not capture the behavior observed in El-Badry et al.
-(2024), and the correction derived in Figure 11 thus only corrects the bias due to to the ruwe  filter.
-We apply our correction to the eccentricity distribution in the NSS astrometric orbits catalog for Sun-like stars (which we roughly define as sources with $64<M_{G,0}<$ 6) and orbital periods $1000100<P_{orb}<$ 1000 (Figure 11) [^11] We note that that these two requirements do not actually change the shape of the eccentricity distribution as compared to the full NSS catalog, they just reduce the number of sources..
-We find that the corrected distribution is very similar to the raw distribution (i.e., in our model, the detection probability for a binary is largely independent of eccentricity), and that the corrected Gaia NSS astrometric orbits catalog still contains significantly fewer high eccentricity binaries than predicted by Moe & Di Stefano (2017).
+Figure 13: Forward model assuming different birth eccentricity distributions.
+The birth distribution of eccentricities for orbital periods within $150030<P_{orb}<$ 1500 days is shown in the dotted blue line.
+The forward modeled eccentricity distribution is shown in the solid blue line.
+The distribution of eccentricities in the NSS astrometric orbits catalog is shown in black.
+The default prescription in Moe & Di Stefano (2017) (left panel) produces an excess of high-eccentricity orbits.
+Assuming a flat eccentricity distribution (left middle panel) $\eta=0$ is a slight improvement.
+Assuming a distribution with $\eta=-0.3$ (right middle panel) matches the long-eccentricity tail well, but overestimates the number of circular orbits.
+A Rayleigh distribution with $\sigma_{e}=0.33$ (right panel) provides the best match to the observed eccentricity distribution.
+As a test of our forward model to see if a different eccentricity distribution would result in a better match to the observed NSS distribution, we re-run the population synthesis, replacing Equation 8 from Moe & Di Stefano (2017) with $\eta=0$ and $-0.3$.
+All other parts of the population synthesis remain the same.
+We then re-run the forward model and look at the resulting eccentricity distribution.
+Figure 13 compares the different eccentricity prescriptions.
+Although neither of them provides a great match, they are improvements at the high eccentricity end.
+The issue with a power law parametrization is that any decrease in high-eccentricity solutions is compensated by more circular solutions.
+Therefore, a different functional form is likely needed.
+The observed NSS eccentricity distribution peaks at $e\approx 0.2-0.3$.
+Qualitatively, the birth distribution needs to have a concentration of solutions around $e=0.2-0.3$ to match the observed Gaia NSS distribution.
+Wu et al. (2025) propose a Rayleigh distribution
+$p(e)=\frac{e}{\sigma_{e}^{2}}\exp\left[-\frac{e^{2}}{2\sigma_{e}^{2}}\right]$ (9)
+to model the eccentricity distribution of Gaia DR3 NSS sources.
+They find that for their “gold sample”, which consists of $\sim 3000$ Sun-like binaries with $10^{2}<P_{orb}<10^{3}$ days within 150 pc, the best-fit scale parameter is $\sigma_{e}=0.336$, which includes a completeness correction to account for NSS selection effects.
+We test this using our forward model.
+For our original synthetic population, for any source with $\log_{10}P_{orb}>0.7$, we instead draw their eccentricities from Equation 9 with $\sigma_{e}=0.33$.
+Because the Rayleigh distribution has support over $[0,\infty)$, we remove any sources that have $e>1$.
+Following Moe & Di Stefano (2017), we also removed sources where the eccentricity exceeded
+$e_{max}(P_{orb})=1-\left(\frac{P_{orb}}{2\textrm{ d}}\right)^{-2/3}$ (10)
+to ensure that the components have Roche lobe fill factors $\lesssim 70\%$ at periastron.
+We simply enforce a hard cutoff at $e_{max}$; in contrast, COSMIC applies a linear turnover in the number of eccentric solutions at $e/e_{max}=0.8$, such that the eccentricity distribution is continuous at $e/e_{max}=0.8$ and zero at $e=e_{max}$.
+Finally, we again run our NSS forward model to determine the resulting eccentricity distribution (rightmost panel of Figure 13), which now closely reproduces the observed NSS distribution.
+As noted in Wu et al.
+(2025), the birth and NSS forward modeled eccentricity distributions are very similar, meaning the eccentricity distribution is approximately complete.
+This is because a Rayleigh distribution with $\sigma_{e}\approx 0.3$ has very few high-eccentricity solutions, where the distribution is most affected by selection effects; it is not generally true of any eccentricity distribution.
 With Gaia DR4 per-epoch astrometry, these findings can be tested and validated.
@@
 Kepler continually monitored a single part of the sky over 4 years and is thus sensitive to eclipsing binaries with orbital periods of up to $\approx 2$ years.
-6.2 Limitations of the approach
-The filters described in §4.2 and 4.3 are empirical, and we do not attempt to derive the relations from first principles nor from the scanning law of the spacecraft.
-The empirical period filters attributed to the loss of orbital solutions assigned to acceleration solutions in the model cascade likely cannot be derived from first principles.
-In addition, it does not include potential correlations with other parameters (e.g., eccentricity) which might affect the probability of a source being given an acceleration solution.
-The empirical period filter is likely adequate for population studies, but less ideal when trying to infer properties of one single system.
-In addition, the empirical period filters do not map directly to a step in the model cascade, as the ruwe  and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters do.
-To do this would require forward modeling individual observations.
-For a more detailed approach that actually does simulate individual Gaia observations including the scanning law, see El-Badry et al.
-(2024) (further discussion is also presented in §6.5).
+6.2 Strengths and limitations of the approach
+In the spirit of keeping calculations fast and analytic in this work, we construct empirical probabilistic functions to approximate the fitting of epoch 1-D astrometry.
+This has both strengths and limitations, namely the tradeoff between speed and simplicity vs.
+computational expense and accuracy.
+The main strength of a fast simulation is for exploring parameter space.
+It requires approximately $10-20$ seconds to run the full astrometric cascade of El-Badry et al.
+(2024) for each source that receives an orbital solution, or about 5000 CPU hours to produce a full mock DR3 binary catalog.
+In contrast, it requires approximately 2 seconds to run 100,000 evaluations of the empirical forward model of this work at a fixed position on sky; most of that time comes from querying the 3-D dust map.
+Of course, it is important to note that the empirical corrections to the analytic prescription are dependent on having the expensive epoch modeling simulations already in hand.
+It is analogous to significant upfront costs in training machine learning models, but later having substantially lower costs when using the model to make predictions.
+Finally, the simple approach also allows for a more intuitive understanding of each step in the cascade and elucidates the complex selection function.
+The speed and simplification comes at the expense of detail and potentially accuracy.
+The major limitation of our analytic models is that they are limited to certain regions of parameter space.
+Most notably, as discussed earlier, our model does not cover orbital periods $P_{orb}>1500$ days.
+In addition, when trying to understand one specific system in detail, individual epoch modeling will likely be more accurate and can provide specific insight.
+One way to take advantage of the strengths of both the fast and detailed approaches are to use the fast simulations to identify potentially interesting regions of parameter space, then to use the detailed simulations to validate and refine those results.
 6.3 Spectroscopic solutions
@@
 6.4 Gaia DR4
-As of late 2024, Gaia DR4 is projected to be released in mid-2026.
+As of 2025, Gaia DR4 is projected to be released in late 2026.
 DR4 will nearly double the time baseline of DR3 (66 months of data vs.
@@
 It is beyond the scope of this work to extend this forward model framework to forecast for Gaia DR4.
-Several parts of this forward model, such as the empirical removal of solutions at certain orbital periods (§4.3), would require knowledge of the DR4 processing pipeline to forecast.
-An alternative way to create a conservative forecast of the selection of DR4 binaries is to model per-epoch photometry and assume that the processing cascade remains the same as that of DR3 (El-Badry et al., 2024).
-However, once DR4 is released, it would not be difficult to extend the model presented here.
+In principle, this can by done by re-generating the catalogs of El-Badry et al.
+(2024) using the 66 month DR4 baseline, and re-calibrating the empirical corrections here.
 Finally, we also note that DR4 will be the first Gaia data release to include individual epoch photometry and astrometry.
 Thus, the completeness and selection function can be evaluated in other manners, such as injection and recovery tests.
-6.5 Comparison to El-Badry et al. (2024)
-In order to characterize the selection function of the Gaia DR3 NSS astrometric orbits catalog, El-Badry et al.
-(2024) forward model individual epoch observations and validate it using a binary population synthesis model.
-Many of the underlying components of the population synthesis are similar or identical to that described in this paper; they also use galaxia and COSMIC to simulate the population of binary stars within 2 kpc.
-Minor differences include the extinction ($A_{G}=2.3E(B-V)$ in this work vs.
-$A_{G}=2.8E(B-V)$ in El-Badry et al.
-2024), number of binaries at birth within 2 kpc (170 million in this work vs.
-195 million in El-Badry et al.
-2024), and the white dwarf initial-final mass relation (Kalirai et al.
-2008 in this work vs.
-Weidemann 2000 in El-Badry et al.
-2024).
-To generate the individual observations, they use GOST to obtain the scan times and scan angles, and use Holl et al.
-(2023) for the noise model of the observations.
-They then fit these mock observations following the cascade described in Halbwachs et al.
-(2023).
-The advantage of modeling the observations at the epoch level is that each step described in the processing pipeline can be modeled.
-The main drawback is the computational expense.
-It requires approximately $10-20$ seconds to run the full astrometric cascade of El-Badry et al.
-(2024) for each source that receives an orbital solution, or about 5000 CPU hours to produce a full mock DR3 binary catalog.
-In contrast, it requires approximately 2 seconds to run 100,000 evaluations of the empirical forward model of this work at a fixed position on sky; most of that time comes from querying the 3-D dust map.
-We also compare our results for the inclusion probability of Gaia BH1, BH2, and BH3 in the Gaia NSS astrometric orbits catalog.
-The results agree well for Gaia BH1 (33% in this work, and 39% in El-Badry et al.
-(2024)).
-For Gaia BH2 and BH3, El-Badry et al.
-(2024) find significantly higher inclusion probabilities (48% and 10%, as compared to the 4% and 0% in this work).
-One difference is that the values for El-Badry et al.
-are sky-averaged, while the values presented here are a function of (RA, Dec).
-For Gaia BH1, the detection probability would be 2×2$\times$2 × higher at higher ecliptic latitudes (Figure 9); for Gaia BH2, the detection probability does not vary across the sky (Figure 9).
-El-Badry et al.
-(2024) also calculate the change in inclusion probabilities as a function of distance.
-For Gaia BH1, the results are qualitatively similar, although our inclusion probability rises more steeply toward closer distances.
-For Gaia BH2, the results disagree, mainly because our inclusion probability does not exceed $\approx 4\%$.
-This may be due to strong covariances between $P_{orb}$, the mass of the dark companion, and detection probability, which is not captured in the probabilistic model that removes long orbital period solutions.
 7 Conclusions
@@
 However, the selection function of this catalog is not well-characterized, and without knowledge of the selection function, discoveries of interesting systems are limited to individual curiosities, as their population statistics cannot be inferred.
-To address this, we create a framework to forward model the selection function of the Gaia DR3 NSS astrometric orbits catalog.
+To address this, we create a fast, simple, and physically interpretable framework to forward model the selection function of the Gaia DR3 NSS astrometric orbits catalog.
+Our method combines analytic prescriptions with detailed results from El-Badry et al.
+(2024).
 Using a binary population synthesis model, we validate this framework.
 We find the method produces reasonable agreement with the observed distribution of periods, parallaxes, photocenter sizes, ruwe, magnitudes, and parallax errors; the one parameter that is substantially discrepant is the eccentricity; we suspect the assumed birth population of eccentricities is the culprit.
-As an example application of this framework, we use it the model to calculate the probability of Gaia BH1, BH2, and BH3 being included in the NSS astrometric orbits catalog.
-We find that Gaia BH1 has a relatively high chance of being included, but Gaia BH2 was a much luckier detection.
-Gaia BH3 was unlikely to be detected, due to its very long orbital period.
+We find the proposed Rayleigh distribution of Wu et al.
+(2025) to better explain the eccentricity distribution of Gaia sources.
+As an example application of this framework, we use it the model to calculate the probability of Gaia BH1 and BH2 being included in the NSS astrometric orbits catalog.
+We find that both Gaia BH1 and BH2 had relatively high probabilities ($>20$%) of being included in the Gaia DR3 NSS astrometric orbits catalog, and that there are a few thousand Sun-like star + BH binaries in the entire Galaxy.
 Ultimately, we look forward to Gaia DR4, where the time baseline will be doubled, and per-epoch astrometry will be released for all sources.
@@
 This work has made use of NASA’s Astrophysics Data System.
-Appendix A Parallax and photocenter orbit semimajor axis error model
-Here we investigate the relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$, and $\sigma_{AL}/\sqrt{N_{vis}}$.
-There is a linear relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$ and $\sigma_{AL}/\sqrt{N_{vis}}$.
-Figure 12 shows this relationship, binned by the magnitude of the source and the number of visibility periods used.
-The binned value is indicated by the dot, the error bar is the error on the mean of that bin.
-Each dot represents a visibility period for that magnitude range; they are connected in order of decreasing number of visibility periods.
-The relationship gets noisy towards fainter magnitudes ($G\gtrsim 16$) and hence the line connecting the dot is no longer straight.
-The relationship for $138<G<$ 13 are all similar because $\sigma_{AL}$ is roughly constant in this magnitude range.
-However, it is unclear why going to fainter magnitudes, $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ are smaller for a given $\sigma_{AL}/\sqrt{N_{vis}}$.
-To investigate, we make the corresponding plot in $\sigma_{\varpi}$ vs.
-$\sigma_{AL}/\sqrt{N_{vis}}$ but using a random subset of the gaia source catalog.
-We find a similar relation persists, albeit the discrepancy at fixed $\sigma_{AL}/\sqrt{N_{vis}}$ is not as pronounced.
-We thus suspect that there is a magnitude dependence on $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ that is not necessarily entirely captured in $\sigma_{AL}$.
-Figure 12: $\sigma_{\varpi}$ vs.
-$\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (top panel), $\sigma_{a_{0}}$ vs.
-$\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (middle panel), and $\sigma_{\varpi}$ vs.
-$\sigma_{AL}/\sqrt{N_{vis}}$ for a random subset of gaia_source sources (bottom panel).
-The sources are binned by magnitude and number of visibility periods used; for a given magnitude, the points are connected by a line indicated increasing number of visibility periods.
-The NSS catalogs exhibit a discontinuity between different magnitudes not seen in the gaia_source catalog.
-Appendix B Cross-match to external catalogs
+Appendix A Eccentricity error $\sigma_{e}$
+Table 1 of Halbwachs et al.
+(2023) lists the criteria for acceptance as an orbital solution as $F_{2}<25$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$, and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$.
+As described in §5.3 “Selection of the orbital solutions” of Halbwachs et al.
+(2023), the criteria $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ only remove an additional 7% of solutions after the other three criteria were applied.
+Here, we provide justification to ignore $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ in our forward modeling.
+We analyze all the orbital solutions in the simulations of El-Badry et al.
+(2024) that pass the filters $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ and $a_{0}/\sigma_{a_{0}}>5$.
+We then 1) remove solutions that also pass the filter $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$, 2) remove solutions that pass the $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filter, and 3) remove solutions that pass both $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filters.
+We show the distribution of $P_{orb}$, $e$, and $a_{0}$ for these solutions in Figure 14.
+The effect of excluding the $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filter makes almost no difference on the inferred parameter distributions.
+Figure 14: Distributions of the orbital period $P_{orb}$, eccentricity $e$, and photocenter orbit semimajor axis $a_{0}$ depending on whether the filters on $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ were included.
+Appendix B Goodness-of-fit $F_{2}$
+The $F_{2}>25$ criterion is mainly for removing spurious solutions.
+Because our modeling assumes the true orbital parameters are known and that all systems are binaries (and not, e.g., triple systems), these spurious solutions do not exist in our forward model.
+We validate this using the simulations in El-Badry et al.
+(2024).
+Figure 15 shows period vs.
+eccentricity for sources where an orbital solution is tried, with bad solutions ($F_{2}>25$) in orange points.
+The left panel shows the true period and eccentricity (which are known because they are simulation inputs).
+The right panel are the inferred values performed from fitting simulated Gaia data (c.f.
+the middle panel of Figure 9 of El-Badry et al.
+2024).
+The solutions removed by the $F_{2}>25$ criterion are all artifacts due to incorrect fits and have true orbital periods $P_{orb}>10,000$ days.
+Given that we only model sources with $P_{orb}<1500$ days, we are justified in assuming that $F_{2}<25$ is always satisfied for our sources.
+Figure 15: Spurious solutions as identified in the epoch modeling of El-Badry et al.
+(2024).
+Sources with $F_{2}>25$ (orange) tend to have incorrectly measured periods.
+The inferred periods (right panel) are much shorter than the true periods ($P_{orb}\gtrsim 10,000$ days), which Gaia DR3 is insensitive to.
+Appendix C Acceleration modeling and orbital properties
+Here we show the correlations between the criteria for accepting an acceleration model and their dependence on the true orbital properties.
+Figure 16 demonstrates the relationship between the significances of the 7- and 9-parameter acceleration solutions $s_{7},s_{9}$ and the parallax over errors $\varpi/\sigma_{\varpi}$ of the respective models.
+As described in §2.3.2 and §4.3 of Halbwachs et al.
+(2023), a 9-parameter solution must have $\varpi/\sigma_{\varpi}>2.1s_{9}^{1.05}$, $F_{2}<25$, and $s_{9}>12$ to be provisionally accepted in the main processing; a 7-parameter solution must have $\varpi/\sigma_{\varpi}>1.2s_{7}^{1.05}$, $F_{2}<25$, and $s_{7}>12$.
+Note that $\varpi/\sigma_{\varpi}$ is different for each different model (7-parameter acceleration, 9-parameter acceleration, orbital model), as different values are inferred when being fit with different NSS models.
+Figure 17 shows the probability of a source to be accepted into the 9-parameter and 7-parameter acceleration model branch of the cascade, as a function of the orbital period and photocenter orbit semimajor axis size.
+The left panel of Figure 17 shows the probability that a source with $\texttt{ruwe}>1.4$ is a 9-parameter acceleration solution candidate as a function of $P_{orb}$ and $a_{0}/\sigma_{AL}$.
+The right panel of Figure 17 shows the probability that a source that has $\texttt{ruwe}>1.4$ and that is not a candidate for a 9-parameter acceleration solution is a candidate for a 7-parameter acceleration solution.
+In both cases, longer period solutions $P_{orb}\gtrsim 500$ days are more likely to have an accepted acceleration solution.
+For 9-parameter acceleration solutions, intermediate $a_{0}/\sigma_{AL}$ sources are preferentially accepted, while for the 7-parameter solutions, lower $a_{0}/\sigma_{AL}$ are preferentially accepted.
+Figure 16: Top four panels: Significance of the 9-parameter acceleration solution (left column) and parallax over error (right column), as a function of true orbital period vs.
+eccentricity (top row), and orbital period vs.
+photocenter orbit semimajor axis (bottom row), for fixed noise (i.e., magnitude and visibility periods).
+Bottom four panels: Same as the top four panels, but for the 7-parameter solutions.
+For both the 9-parameter and 7-parameter acceleration solutions, the significance and parallax over error are strong functions of the orbital period and the size of the photocenter orbit semimajor axis, but do not depend on the orbital eccentricity.
+Figure 17: Probability of acceptance of the 9-parameter (left) and 7-parameter acceleration solutions (right), as a function of the true orbital period and the photocenter orbit semimajor axis, scaled by the per-observation uncertainty $\sigma_{AL}$ (which is a function of magnitude).
+The probability of 9-parameter solution acceptance is relative to all sources that have $\texttt{ruwe}>1.4$; the probability of 7-parameter solution acceptance is relative to all sources that did not have a 9-parameter solution accepted.
+Appendix D Cross-match to external catalogs
 To test the hypothesis that the decrease in solutions with orbital periods $P_{orb}>550$ days is due to the model cascade and losing solutions to acceleration solutions, we cross-matched the Gaia DR3 NSS astrometric orbits and acceleration catalogs to two other binary star catalogs.
-B.1 Cross-match to SB9
+D.1 Cross-match to SB9
 First, we cross-matched the SB9 catalog (Pourbaix et al., 2004) to the Gaia DR3 nss_acceleration_astro  and nss_two_body_orbit  catalogs to empirically determine the number of orbital solutions lost to the acceleration catalog.
@@
 However, this test suggests that our hypothesis that the decrease in solutions with $1000550<P_{orb}<$ 1000 days is due to the model cascade is plausible.
-B.2 Cross-match to WDS-ORB6
+D.2 Cross-match to WDS-ORB6
 Next, we perform the same cross-matching exercise instead using the Sixth Catalog of Orbits of Visual Binary Stars (Orb6, an updated version of Hartkopf et al.